### v2

# Incremental Generation is Necessary and Sufficient for Universality in Flow-Based Modelling

Abstract
Incremental flow-based denoising models have reshaped generative modelling, but their empirical advantage still lacks a rigorous approximation-theoretic foundation. We show that incremental generation is necessary and sufficient for universal flow-based generation on the largest natural class of self-maps of $0,1]^{d}$ compatible with denoising pipelines, namely the orientation-preserving homeomorphisms of $0,1]^{d}$. All our guarantees are uniform on the underlying maps and hence imply approximation both samplewise and in distribution.
Using a new topological-dynamical argument, we first prove an impossibility theorem: the class of all single-step autonomous flows, independently of the architecture, width, depth, or Lipschitz activation of the underlying neural network, is not universal in the space of orientation-preserving homeomorphisms of $0,1]^{d}$. By exploiting algebraic properties of autonomous flows, we conversely show that every orientation-preserving Lipschitz homeomorphism on $0,1]^{d}$ can be approximated at rate $\mathcal{O}(n^{-1/d})$ by a composition of at most $T_{d}$ such flows, where $T_{d}$ depends only on the dimension. Under additional smoothness assumptions, the approximation rate can be made dimension-free, and $T_{d}$ can be chosen uniformly over the class being approximated. Finally, by linearly lifting the domain into one higher dimension, we obtain structured universal approximation results for continuous functions and for probability measures on $0,1]^{d}$, the latter realized as pushforwards of empirical measures with vanishing $1$-Wasserstein error.
Keywords: Incremental Generation; Normalizing Flows; Neural ODEs; Universal Approximation; Dynamical Systems; Homeomorphism Groups; Flowability; Optimal Transport, Structure-Aware AI.
MSC (2020): 41A30, 41A65, 58D05, 58D15, 37C10.
1 Introduction
Flow-based diffusion models are changing our understanding of generative modelling, combining exact likelihoods with high-quality samples rezende2015variational; kingma2018glow, and the effectiveness of the modern incremental generation paradigm has undeniably revolutionized a range of areas from image synthesis to molecular design and time-series modelling papamakarios2021normalizing; ho2020denoising; kobyzev2020normalizing. Nevertheless, our fundamental question remains: why and how is incremental generation advantageous over (non-incremental) predecessors such as GANs goodfellow2014generative. We answer this question in the foundations of AI, using an approximation-theoretic lens.
Here, a generative model refers to any function that transforms a source of noise, given by a random variable $Z$ on $\mathbb{R}^{d}$, into another “generated” random variable
$X=\varphi(Z),$ (1)
where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{D}$. Generation is performed by sampling $Z$, which induces a sample of $X$. The transformation $\varphi$ is required to be continuous, often Lipschitz, to ensure numerical stability during generation gulrajani2017improved; cisse2017parseval; miyato2018spectral. We consider denoising-based generative models, such as normalizing flows rezende2015variational; kingma2018glow which conceptually build on their variational autoencoder (VAE) predecessors kingma2013auto. These models require the transformation $\varphi$ to be continuously invertible and leverage the inverse map $\varphi^{-1}:\varphi(\mathbb{R}^{d})\to\mathbb{R}^{d}$ to train the model: given i.i.d. samples from $X$, the model learns to map them back to the noise source $Z$. The invertibility, thus injectivity of $\varphi$ necessitates that $D\geq d$; we focus on the difficult critical case where $D=d$ and $\varphi$ is surjective (other cases are a simple consequence which we discussed below). Lastly, keeping with the approximation-theoretic tradition in deep learning mhaskar2016deep; yarotsky2017error; petersen2018optimal; elbrachter2021deep; kratsios2022do; zhang2022deep; siegel2023optimal, we will focus only on maps $\varphi$ that are supported (in the sense of differential topology, cf. (Banyaga_1997__DiffeoBook, Chapter 2)) on the unit hypercube $0,1]^{d}$; that is, $\varphi(x)=x$ for every $x\notin[0,1]^{d}$. Consequently, Alexander’s trick implies that every such homeomorphism must be orientation preserving. This means that, the largest concept class of denoising-type generative models consists of all random variables $X$, as in (1), for which the transformation $\varphi$ belongs to the class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ of orientation-preserving homeomorphism supported on $0,1]^{d}$. Whenever both $X$ and $Z$ have a Lebesgue density on $0,1]^{d}$ then the existence of such a $\varphi$ is guaranteed by the main result of dacorogna1990partial; regardless of $Z$ and $X$.
[…]
$\sup_{\omega}\,\|\varphi(Z(\omega))-\psi(Z(\omega))\|\leq d_{\infty}(\varphi,\psi)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\sup_{x\in\mathbb{R}^{d}}\,\|\varphi(x)-\psi(x)\|$ (2)
where the supremum on the left-hand side of (2) is taken over all outcomes $\omega$ in the probability space on which $Z$ is defined. We emphasize that $d_{\infty}(\phi,\psi)$ must be finite since both $\phi$ and $\psi$ coincide outside of $0,1]^{d}$. In other words, the relevant concept class in this paper is the function space $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ with distance quantified by $d_{\infty}$, which will serve as the focus of all our approximation-theoretic analysis.
Unfortunately, $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is not a vector space which places it outside the scope of the available constructive approximation toolbox pinkus2012n; CABook2; cohen2022optimal; petrova2023Lipschitz, which assumes a Banach space structure—our concept class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ possesses a rich and well-studied infinite-dimensional topological group structure mather1974commutators; thurston1974foliations; fathi1980structure; mather1984curious; banyaga1997structure; haller2013smooth; fukui2019uniform that we will exploit extensively. Moreover, unlike signature-based machine learning methods, cf. gyurko2013extractingsignature; chevyrev2016primer; cuchiero2023signaturecalibration; hamida2005signaturetest; cass2024lecturenotes; bayer2025optimalstopping, rooted in rough path theory lyons1998differential; gubinelli2004controlling; HamblyLyons2010, we do not have access to a global exponential map (in the sense of infinite-dimensional Lie groups; cf. Kriegl1997; michor2023regularity) to transfer results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$, as in kratsios2020non; kratsios2022universalpapon. In short, there are currently no available tools to apply classical approximation results from linear spaces onto $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ using global exponential map. we are required to develop new approximation-theoretic techniques suited to its non-vectorial structure, where our approximators/hypothesis class consists only of (neural network-based) homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. Our approximators preserve the structure of the maps in (1), whereas replacing $\varphi$ with an arbitrary, e.g. non-invertible neural network as in the standard GAN framework radford2016unsupervised; arjovsky2017wasserstein; peyre2019computational; rout2022generative; korotin2021wasserstein; korotin2023neural; kolesov2024energy prevents the use of denoising-based training.
Semi-Discrete Dynamics
We identify and study the approximation-theoretic gap between incremental and non-incremental flow-based generative models. In this paper, a flow-based generative model means a map $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ decomposable as a ”semi-discrete dynamical system”; by which we mean $\varphi$ can be expressed as the composition of finitely many (discrete part) flows of continuous-time ODEs (continuous part). That is
$\varphi$ $=\varphi_{T}\circ\dots\circ\varphi_{1},$ (3)
for some $T\in\mathbb{N}_{+}$, where for each $l\in\{1,\dots,T\}$, the map $\varphi_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a time-$1$ flow of an autonomous ODE; that is, $\varphi_{l}$ sends any $x\in\mathbb{R}^{d}$ to the time-$1$ solution $x_{1}^{x}$ of the autonomous ODE with initial condition $x$
$\frac{dx_{t}^{x}}{dt}$ $=V_{l}(x_{t}^{x}),\quad x_{0}^{x}=x$ (4)
where $V_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a Lipschitz vector field, parameterized by a ReLU multilayer perceptron (MLP)whose weights matrices has finite operator norm. Under this formalism, a flow-based generative model $\varphi$, as in (3), is called incremental if $T>1$ and it is called non-incremental if $T=1$. We note some $L^{p}$-version, for finite $p$, of these results in this direction are known via controllability arguments ruiz2024control; scagliotti2023deep; cuchiero2020deep; cheng2025interpolation, however these results come at a cost as they require the user to have active control of the vector field of the Neural ODE (i.e. non-autonomous or controlled Neural ODEs) which is effectively the case where $T=\infty$.
Figure 1: Visualizing Incremental Flow-Based Generation (3): In a denoising-type, flow-based incremental generator, an image $x$ (left) is mapped to noise by the invertible sequence $\varphi_{T}^{-1}\circ\cdots\circ\varphi_{1}^{-1}$. The model is trained to undo this via $\varphi_{1}\circ\cdots\circ\varphi_{T}$. At inference, a noise sample $Z$ is injected and propagated through $\varphi_{1},\dots,\varphi_{T}$ to synthesize an image  [^2] Image credit: Sans Soleil (1983), written and directed by Chris Marker..
Here, $T=2$. Each arrow depicts the vector field—visualized in the panel directly below—that induces one step of the incremental flow. By contrast, a non-incremental generator attempts a single-shot mapping from left to right, while non-denoising pipelines (e.g., GANs goodfellow2014generative) do not enforce invertibility.
In contrasts, incremental flow-based generative models can be approximately represented as a single non-autonomous flow, i.e., where the vector field in (4) is allowed to be time-dependent but undergoes only finitely many changes in direction; which directly allows the conversion of our theory of incremental flow-based generation to time-dependent generation. Such time-dependent models differ from controlled neural ODEs kidger2020neural; cuchiero2020deep; cirone2023neural, which mimic controlled rough paths morrill2021neural; pmlr-v235-walker24a and can potentially adapt their vector fields arbitrarily often over time resulting in a higher parametric complexity. The following question drives our manuscript and an affirmative answer to it would imply a concrete mathematical support for the advantage of incremental generation over classical non-incremental generation to support the undeniable success of incremental generation
Question 1 Is there any approximation theoretic advantage of incremental generation?
1.1 Main Construction
Our first result (Theorem 2) shows that incremental generation, i.e. when $T>1$, is necessary for universality of flow-based generative models in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for $d\in\mathbb{N}_{+}$ and $d>1$. Specifically, there exists a $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ that cannot be approximated by the flow of any autonomous neural ODE, regardless of network depth, width, or choice of activation function. We deduce this from a more general result of ours showing that: the class of all flows of autonomous ODEs is non-universal (in particular, not dense) in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$. This result extends the differential topology results of Varandas_2020__Meag to the non-compact case and the classical embedding results of fort1955embedding; jones1972embedding; palis1974vector to the non-smooth, non-compact setting.
Our second result (Theorem 8) shows that incremental generation with ReLU vector fields is sufficient. Moreover, the result is quantitative, and, surprisingly, we find that the added homeomorphism structure allows our approximation rates to match the minimax-optimal rates achieved by ReLU neural networks in the broader classes of uniformly continuous yarotsky2017error; shen2020deep and $C^{k}$-functions petersen2018optimal; yarotsky2020phase; lu2021deep from $0,1]^{d}$ to $\mathbb{R}^{d}$, respectively.
Our final main result (Theorem 10) shows, somewhat surprisingly, that $T$ never needs to become arbitrarily large when approximating diffeomorphisms. Even more strikingly, we prove that there exists a dimensional constant $T_{d}>0$ such that one can constrain $T\leq T_{d}$, independently of the target diffeomorphism being approximated and maintain universality in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
1.2 Secondary Contributions
2 Preliminaries
2.1 Background
2.1.1 Homeomorphisms, Flows and Flowability
This section contains all necessary terminology and background needed to formulate our main results. Additional background required only for proofs is contained in Appendix D. We say a function is supported on a set $S$ if it is zero out of that set and a function is supported if it is the identity out of that set. For practical reasons said in the introduction section, we need to work with compactly supported/supported functions and vector fields; meaning that $S$ is contained in a compact set.
We recall that, a homeomorphism is a continuous bijective function with continuous inverse. Furthermore, if the function and its inverse are $C^{k}$-differentiable, it is called $C^{k}$-diffeomorphism. By convention, $C^{\infty}$-diffeomorphism is called diffeomorphism. Two continuous functions $f,g:X\to Y$ between the same topological spaces $X$ and $Y$ are thought of as being “topologically similar” if one can be continuously deformed into another; such as continuous deformation is called a homotopy; i.e. $H(t,x):[0,1]\times X\to Y$ such that $H(0,\cdot)=f$ and $H(1,\cdot)=g$. We will say that two homeomorphisms are isotopic if one can continuously transform one into the other with homeomorphisms.
Example 1
If $d,D\in\mathbb{N}_{+}$ and $f,g:\mathbb{R}^{d}\to\mathbb{R}^{D}$ are continuous then the so-called linear homotopy $H:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{D}$ sending any pair $(t,x)$ to (1t)f(x)+tg(x)(1-t)f(x)+tg(x) is indeed a homotopy.
[…]
Example 2 (A $1d$ Example of (Non-)Orientation Preserving Homeomorphism)
A transparent example of an orientation preserving homeomorphism in one dimension is the identity $\mathbb{R}\ni x\to x\in\mathbb{R}$. A homeomorphism from $\mathbb{R}$ to itself which is not is the mirroring map $\mathbb{R}\ni x\mapsto-x\in\mathbb{R}$. Indeed, both of these maps are homeomorphisms since they are their own inverses.
Example 3 (Orientation-Preserving Homeomorphisms From Computer Vision (Rotations))
Standard Multidimensional examples arising in rotation invariances in computer vision, e.g. lui2012human; cohen2016group; thomas2018tensor, include linear transformerization $\varphi:\mathbb{R}^{d}\ni x\mapsto Ox\in\mathbb{R}^{d}$ where $O$ is an $d\times d$ orthogonal matrix; in which case $\varphi$ is orientation preserving if and only if $\det(O)=0$ and it fails to be precisely when $\det(O)=-1$.
If $O$ is orientation-preserving; i.e. if $\det(O)=1$, then any such homeomorphism can be expressed as the solution to an ordinary differential equation (ODE) at time $1$; namely, $\varphi(x)=Ox=x_{1}^{x}$ where
$\tfrac{d}{dt}\,x_{t}^{x}$ $=\mathfrak{o}x_{t}^{x}$ (5)
$x_{0}^{x}$ $=x$
where $O=\exp(\mathfrak{o})$ for some $d\times d$-skew-symmetric matrix $\mathfrak{o}$; where $\exp$ is the matrix exponential. This can be noted upon observing that the ODE solution to the ODE (5) is given by the curve $x_{\cdot}^{x}=(x_{t}^{x})_{t\geq 0}$ where
$x_{t}^{x}=\exp(t\mathfrak{o})\,x.$
This connection is the starting-point of the theory of finite-dimensional Lie groups; cf. helgason1979differential. Indeed, finite-dimensional Lie groups induce the prototypical and simplest classes of “well-behaved” homeomorphisms on $\mathbb{R}^{d}$ in this way; however, only the identity is compactly supported in any such construction where the vector field in (5) is “constant” multiplication against a single matrix.
Generalizing Example 3 we bring our attention to autonomous ODEs of the form (4), where vector-field $V:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is allowed to be any Lipschitz or $C^{k}$-differentiable vector field compactly supported on interval domains (i.e. $a,b]^{d}$ for $a,b\in\mathbb{R}$). By the Picard-Lindelöf Theorem (Hartman_2002__ODE, Chapter II Theorem 1.1) this system has a unique global solution $x_{t}^{x}$ and the map sending the point $x$ to the time-1 value $x_{1}^{x}$ of the solution to this problem (often referred to as Cauchy problem) will be called the flow induced by the vector field $V$, and is denoted by $\operatorname{Flow}{}(V)$. The flow of any Lipschitz vector field supported on a compact set $S$ defines a compactly supported homeomorphism, supported on $S$. If, moreover, $V$ is $C^{k}$-differentiable then its flow is actually a $C^{k}$-diffeomorphism. We say that an orientation-preserving homeomorphism in $d$ dimensional Euclidean space, $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is flowable if there exists a Lipschitz Vector field $V$ supported within $0,1]^{d}$ such that $\varphi=\operatorname{Flow}{}(V)$. An interesting connection, worth noting here in analogy with Example 3—is that the spaces of flows and diffeomorphisms Banyaga_1997__DiffeoBook constitute the prototypical infinite-dimensional Lie groups, a fact that sparked the foundational work of kriegl1997regular; kriegl1997convenient; teichmann2001regularity; neeb2007infinite; omori2017infinite; michor2023regularity.
2.1.2 From Neural ODEs to Incremental Flow-Based Generators
Let $\Delta\in\mathbb{N}_{+}$ and consider a multi-index $\mathbf{d}\ \stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\ [d_{1},\ldots,d_{\Delta+1}]^{\top}\in\mathbb{N}_{+}^{\Delta+1}$. The class $\mathcal{NN}(\mathbf{d})$ consists of all multilayer perceptrons with $\sigma\in C(\mathbb{R})$ Lipschitz-activation function ($\sigma$-MLPs)
$\Phi:\ \mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{\Delta+1}}$
admitting the following iterative representation
$\Phi(\mathbf{x})$ $=\mathbf{W}^{(\Delta)}\mathbf{x}^{(\Delta)}+\mathbf{b}^{(\Delta)},$ (3.2)
$\mathbf{x}^{(l+1)}$ $\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\ \sigma\bullet\!\bigl(\mathbf{W}^{(l)}\mathbf{x}^{(l)}+\mathbf{b}^{(l)}\bigr),$ $\text{for }l=1,\ldots,\Delta-1,$
$\mathbf{x}^{(1)}$ $\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\ \mathbf{x}\,.$
Here, for $l=1,\ldots,\Delta$, $\mathbf{W}^{(l)}$ is a $d_{l+1}\times d_{l}$ matrix and $\mathbf{b}^{(l)}\in\mathbb{R}^{d_{l+1}}$, and $\sigma\ \bullet$ denotes componentwise application of the $\sigma$ function. Fix a Lipschitz activation function $\sigma\in C(\mathbb{R})$ and real numbers $a\leq b$. A $\sigma$-neural ODE of respective depth and width $\Delta,W\in\mathbb{N}_{+}$ is a flowable homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ for which there exists a $\sigma$-MLP $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ of depth $\Delta$ and width $W$ supported in $a,b]^{d}$ such that $\varphi$ is time one solution of (4) with vector field $\Phi$; also, the class of all $\sigma$-neural ODEs supported on a set $P\subseteq\mathbb{R}^{d}$ is denoted as $\mathrm{NODE}_{\sigma}(P)$. The class of incremental flow-based generators with activation function $\sigma$ denoted as $\mathrm{IFG}_{\sigma}(P)$ is defined as all functions of $\operatorname{\mathcal{H}}_{d}(P)$ that can be written as composition of at least two or more but finitely many $\sigma$-neural ODEs.
2.2 Notation
Before moving on, we now collect a list of notation used in our manuscript.
  • Let $\mathbb{N}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{0,1,2,\dots,\}$ , $\mathbb{N}_{+}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{n\in\mathbb{N}:\,n>0\}$ and for $N\in\mathbb{N}_{+}$ denote $N]\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{1,\dots,N\}$
  • For a Lipschitz function $f$ we denote its Lipschitz constant as $L^{f}$
  • Given a function $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$, we denote its support by $\mathrm{supp}(f)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{x\in\mathbb{R}^{d}:\,f(x)\neq 0\}$ and its support by $\mathrm{supp}^{\dagger}(f)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\overline{\{x\in\mathbb{R}^{d}:\,f(x)\neq x\}}=\overline{\mathrm{supp}(f-\mathrm{id})}$
  • For a set $P$ denote its interior by $int(P)$
  • Denote (open Euclidean) ball with radius r around point p as $B_{r}(p)=\{x\in\mathbb{R}^{d}|\ \|x-p\|<r\}$ where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
  • Let $\operatorname{\mathcal{H}}_{d}$ be the set of orientation preserving homeomorphisms of $\mathbb{R}^{d}$. Given subset $P\subset\mathbb{R}^{d}$, let $\operatorname{\mathcal{H}}_{d}(P)$ be the set of orientation preserving homeomorphisms of $\mathbb{R}^{d}$ compactly supported on $P$
  • For $s\in\mathbb{N}\cup\{\infty\}$ let $\mathcal{X}^{s}$ be the set of $C^{s}$ vector fields on $\mathbb{R}^{d}$ supported on $0,1]^{d}$. Moreover, we consider $\mathcal{X}^{0}$ to be the set of Lipschitz continuous vector fields.
  • For a given vector field $V\in\mathcal{X}^{0}$ let $\mathrm{Flow}(V)$ be the time one solution of:
    $\begin{split}\begin{cases}\dot{x}_{t}^{x}&=V(x_{t}^{x})\\ x_{0}^{x}&=x.\end{cases}\end{split}$ (6)
    and $\mathrm{Flow}([a,b]^{d})$ be the set of time one solutions above with vector fields supported on $a,b]^{d}$ and
    $\mathrm{Flow}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcup_{\forall k\in\mathbb{N}\cup\{\infty\}}\{\mathrm{Flow}(V)|V\in\mathcal{X}^{k}\}$
    also $\mathrm{Flow}_{\sigma}([a,b]^{d})$ be the set of time one solutions with $\sigma$ neural vector fields supported on $a,b]^{d}$
  • Fix an Incrementality $T\in\mathbb{N}_{+}$ and a smoothness $s\in\mathbb{N}$. A homeomorphism $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ belongs to $\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ if and only if:
    1. (i)
      Representation: There exist vector fields $V_{1},\dots,V_{T}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{t=1}^{T}\,\operatorname{Flow}{}(V_{t}).$ (7)
    2. (ii)
      Minimality: There is no integer $1\leq\tilde{T}<T$ and vector fields $\tilde{V}_{1},\dots,\tilde{V}_{T}\in\mathcal{X}^{s}$ such that
      $\varphi=\bigcirc_{t=1}^{\tilde{T}}\,\operatorname{Flow}{}(\tilde{V}_{t}).$
  • Let $\mathrm{Diff}_{0}(P)$ be the set of orientation preserving diffeomorphisms of $\mathbb{R}^{d}$ compactly supported on the set $P$ .
[…]
  • Given $E\subseteq\mathbb{R}^{d}$, $d$-tuple $\alpha=[\alpha_{1},\alpha_{2},\cdots,\alpha_{d}]^{\top}\in\mathbb{N}^{d}$ and functions $f(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ and $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ let :
    • $\|f(x_{0})\|_{l^{\infty}}=\|(f_{1}(x_{0}),\dots,f_{D}(x_{0}))\|_{l^{\infty}}=\text{max}_{i\in[D]}f_{i}(x_{0})$
    • $\|g\|_{L^{\infty}(E)}=\text{ess sup}_{x\in E}\,|g(x)|$
    • $\|f\|_{L^{\infty}(E)l^{\infty}}=\|\text{ess sup}_{x\in E}\,|f_{1}(x)|\,,\dots,\,\text{ess sup}_{x\in E}\,|f_{1}(x)|\,\|_{l^{\infty}}=\text{max}_{i\in[D]}\,\text{ess sup}_{x\in E}\,f_{i}(x_{0})$
    • $\|\alpha\|_{1}=|\alpha_{1}|+|\alpha_{2}|+\cdots+|\alpha_{d}|$
    • $\partial^{\alpha}=\frac{\partial^{\alpha_{1}}}{\partial x_{1}^{\alpha_{1}}}\frac{\partial^{\alpha_{2}}}{\partial x_{2}^{\alpha_{2}}}\cdots\frac{\partial^{\alpha_{d}}}{\partial x_{d}^{\alpha_{d}}}$
    • $\|g\|_{C^{s}(E)}:=\max\left\{\|\partial^{\alpha}g\|_{L^{\infty}(E)}:\alpha\in\mathbb{N}^{d}\ \text{with}\ \|\alpha\|_{1}\leq s\right\}$ [^4] Note that $\|g\|_{C^{0}}=\|g\|_{L^{\infty}}$
  • Compositional Notation: During the course of our analysis, it will be convenient to describe ReLU MLPs via the role of each of their (sets of) layers. Specifically, the structure of a ReLU MLP $\Phi$ is represented in the following way: suppose $\Phi\;=\;\mathcal{L}_{m}\circ(\sigma\circ\mathcal{L}_{m-1})\circ\cdots\circ(\sigma\circ\mathcal{L}_{2})\circ(\sigma\circ\mathcal{L}_{1})$ where the $\mathcal{L}_{i}$’s are affine transformations. Following Hong_2024__ReLUMLPs, we express this notationally as
    $\mathbf{x}$ $\Longrightarrow(\sigma\!\circ\!\mathcal{L}_{1})(\mathbf{x})\Longrightarrow(\sigma\!\circ\!\mathcal{L}_{2})\circ(\sigma\!\circ\!\mathcal{L}_{1})(\mathbf{x})$
    $\Longrightarrow(\sigma\!\circ\!\mathcal{L}_{m-1})\circ\cdots\circ(\sigma\!\circ\!\mathcal{L}_{2})\circ(\sigma\!\circ\!\mathcal{L}_{1})(\mathbf{x})$
    $\Longrightarrow\mathcal{L}_{m}\circ(\sigma\!\circ\!\mathcal{L}_{m-1})\circ\cdots\circ(\sigma\!\circ\!\mathcal{L}_{2})\circ(\sigma\!\circ\!\mathcal{L}_{1})(\mathbf{x})$
    $=\Phi(\mathbf{x})\,.$
    In other words, if $x_{1},x_{2},\ldots,x_{m-1}$ are the $1,2,\ldots,(m-1)$-th hidden layers of $\Phi$ and $x_{m}$ is the output layer, then the structure of $\Phi$ is expressed as
    $x\;\Longrightarrow\;x_{1}\;\Longrightarrow\;x_{2}\;\Longrightarrow\;\cdots\;\Longrightarrow\;x_{m-1}\;\Longrightarrow\;x_{m}\;=\;\Phi(x)\,.$
3 Main Results
3.1 Negative Results: Incremental Generation is Necessary
Our first main result shows that the collection of all flowable homeomorphisms compactly supported on $0,1]^{d}$ is non-universal in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$; moreover, the collection of all flowable homeomorphisms compactly supported on $\delta,1-\delta]^{d}$ for any given $0<\delta<\tfrac{1}{2}$ are a small subset of homeomorphisms compactly supported on $0,1]^{d}$. This means there is not only one but many functions in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ which cannot be approximated by flowable functions in $\operatorname{\mathcal{H}}_{d}([\delta,1-\delta]^{d})$ for any $0<\delta<\tfrac{1}{2}$.
(a) The orbits of any (topologically) generic dynamical system (yellow) are perturbed to become periodic, by perturbing the underlying system (gray) – $C^{0}$-Closing lemma.
(b) The dynamical system is then further perturbed so that each periodic point becomes a (non-fixed) periodic attractor; which cannot happen for any flow.
Figure 2: Why Non-Incremental Generation is Not Universal: The reason why non-incremental generators (Theorem 2) fail to be universal is that most homeomorphisms cannot be represented as flows (Theorem 3). The idea is that there is a dense open set of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$, which can be approximated/perturbed so that any given orbit becomes periodic (SubFigure 2(a)). Then, these perturbations can be further perturbed so that the a small neighbourhood around the given orbit becomes a basin of attraction (SubFigure 2(b)), which cannot happen for any flow. Consequently, the complement of any such map, which contains the set $\mathrm{NODE}_{\sigma}((0,1)^{d})$ cannot be dense, implying that all non-incremental generators/autonomous Neural ODEs fail to be universal approximators of orientation-preserving homeomorphisms supported on the hypercube $0,1]^{d}$.
Theorem 1 (Incremental Generation is Necessary for Universality)
$\mathrm{NODE}_{\sigma}([0,1]^{d})$ is not universal in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Theorem 2 (Non-Incremental Generation on $\delta,1-\delta]^{d}$ is meagre)
Let $d\in\mathbb{N}_{+}$ with $d>1$ then for any Lipschitz activation function $\mathrm{NODE}_{\sigma}([\delta,1-\delta]^{d})$ is nowhere dense in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Theorem 2 is implied by our more general result in topological dynamics, illustrated in Figure 2,
Theorem 3 (Few $C^{0}$ Homeomorphisms are Flowable on $0,1]^{d}$ for $d>1$)
Let $d\in\mathbb{N}_{+}$ with $d>1$. The set of flowable homeomorphisms in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is meagre in $\operatorname{\mathcal{H}}_{d}(P)$ for every $d>1$ and compact set $P$ s.t. $0,1]^{d}\subset int(P)$.
Remark 4
For simplicity and convention $0,1]^{d}$ is considered but this theorem is also true for any compact set $K,P$ with $K\subset int(P)$. To understand and see the proof refer to Appendix A.
3.2 Positive Results: Incremental Generation with ReLU MLP Vector Fields is Sufficient
3.2.1 Qualitative Universal Approximation Guarantee
We now contrast our previous result, which shows the limitations of non-incremental generation with our universal approximation guarantee, showing that the set of incremental flow-based generators $\mathrm{IFG}_{\sigma}([0,1]^{d})$ is universal in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$; quantitatively. We emphasize our critical structural point that the approximation is “within” the class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ not from “outside”; by which we mean that our approximating class only consists of homeomorphisms not standard continuous functions.
We now state the streamlined qualitative version of our main result, in the high-dimensional setting where $d\geq 5$, before entering into a more technical analysis involving approximation rates and exactly parameter estimates, shortly. The main intuition behind this result is summarized in Figure 3, whose proof technique largely combines both algebraic and approximation theoretic tools.
Theorem 5 (Universal Approximation of Orientation-Preserving Homeomorphisms)
Let $d\in\mathbb{N}_{+}$ and $d\geq 5$. There exists a constant $T_{d}\in\mathbb{N}_{+}$ such that: for every $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ and every $\varepsilon>0$ there exists some $T\leq T_{d}$ and ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that the homeomorphism $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
$\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\varepsilon.$ (8)
(a) Approximation of homeomorphisms using decomposition of diffeomorphisms into composition of flows.
(b) Approximation of each flow by ReLU MLP approximation of its vector field.
Figure 3: By (a uniform version) of Thursten’s Theorem thurston1974foliations we $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ is a simple group and since the group generated by flows is normal, then every diffeomorphism (green curve) must be the composition of finitely many flows of vector fields $V^{(1)}$, $\dots$, $V^{(T)})$ (here $T=4$) – Sub-Figure 3(a). Each vector field is then approximated by a ReLU MLP at an optimal rate (Sub-Figure 3(b)) with maximal Lipschitz regularity using Hong_2024__ReLUMLPs; the approximation of the original homeomorphism (Theorem 8) is concluded using Grönwall’s inequality.
Furthermore, in dimension $d\geq 5$, every orientation preserving homeomorphism is isotopic to the identity, then muller2014uniform implies that it can be uniformly approximated by diffeomorphisms; reducing (Theorem 5) to the smooth case.
Remark 6
For simplicity and convention $0,1]^{d}$ is considered but all of the theorems in this section are also true for any compact set PP.(JMLR:v25:23-0912,yarotsky2018optimal,petersen2018optimal) To understand and see the proof refer to Appendix B.
3.2.2 Quantitative Formulations
We start this section by remarking that, the class $\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ for some $T>1$ is non-empty. Thus the theorem provides a non-vacuous statement for $T>1$.
Proposition 7 (Non-triviality)
there exists some $T\in\mathbb{N}_{+}$ and $T>1$ such that $\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ is non-empty,
We now provide a detailed quantitative analysis of the general qualitative result of the previous section.
Theorem 8 (Universal Approximation by Deep Neural ODEs)
Let $n,d\in\mathbb{N}$, and $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,0}([0,1]^{d})$ then, there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$ (9)
The right hand side converges to zero as $n\rightarrow\infty$.
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLP vector fields of depth $\lceil\text{log}_{2}d\rceil+6$, width $8d(n+1)^{d}+9$, and at-most $16d(n+1)^{d}+9$ non-zero parameters. In particular, these vector fields and $L^{\Psi}$ do not depend on the parameter $n$.
Furthermore, $\omega^{(t)}$ is the modulus of regularity of $\varphi^{(t)}$ which is equal to the modulus of regularity of $\Phi^{(t)}$.
Neural ODE $\Psi$ is compactly supported on $(0,1)^{d}$ showing that incrementality breaks meagreness of $\mathrm{NODE}([\delta,1-\delta]^{d})$ for any $\delta\in(0,\frac{1}{2})$.
Naturally, one may wonder if improved rates are achievable under additional smoothness of the target homeomorphism. Indeed, we confirm that this is the case, mirroring the classical approximation theory of smooth functions by ReLU MLPs.
Proposition 9 (Universal Approximation by Deep Neural ODEs (Differentiable Case))
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ and $N,\,L,\,s\in\mathbb{N}_{+}$ then, there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N,L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\end{split}$ (10)
The right hand side goes to zero as $L\rightarrow\infty$ and $N\rightarrow\infty$.
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}_{+}$. In particular, $L^{\Psi}$ depends on the parameter $N$.
Furthermore, $(\omega^{(g)})_{j}=85(s+1)^{d}8^{s}\|V^{(g)}_{j}\|_{C^{s}([0,1]^{d})}(NL)^{-2s/d}$.
Complexity of MLP Vector fields Lipschitz Case $C^{s}$-Differentiable Case ($s\geq 1$)
Depth $\lceil\log_{2}(d)\rceil+4$ $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$
Width $8d(n+1)^{d}$ $17sd^{\,d+13}d^{2}(N_{j}+2)\log_{2}(8N_{j})$
Nonzero parameters $16d(n+1)^{d}$
Table 1: Parametric Complexity of the Flow-Based Generative Model of Theorem 8 and of Proposition 9.
Now using the theorem bellow we give an order of approximation for any diffeomorphism in Proposition 11.
Theorem 10 (Finite Composition of Flows for Diffeomorphisms)
There exists a positive number $T_{d}\in\mathbb{N}_{+}$ such that any diffeomorphism $\varphi\in\text{Diff}_{0}([0,1]^{d})$ can be written as at most $T_{d}$ flows.
Proposition 11 (Universal Approximation by Deep Neural ODEs (Smooth Case))
Let $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ and $N,L\in N_{+}$ and take $L$ fixed; then, for $\forall s\in\mathbb{N}_{+}$ there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\in O\big(N^{-2s/d}\big)\end{split}$ (11)
Moreover, $\Psi$ is a compactly-supported homeomorphism on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Phi^{(1)},\dots,\Phi^{(T)}$ are ReLU MLPs of width less than or equal to $17s^{d+1}3^{d}d^{2}(N+2)\log_{2}(8N)$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ where $N,L\in\mathbb{N}^{+}$. In particular, $L^{\Psi}$ depends on the parameter $n$.
4 Implications: Lifted Flow-Based Generation Imply Universal Approximation
We now present a sequence of corollaries that further illustrate the scope of our results. In particular, the simplest form of our main positive result—our approximation theorem—yields structured (in the sense of homeomorphisms “lying over,” as explained below) versions of state-of-the-art universal approximations for ReLU MLPs, implemented via a “tweaked” incremental flow-based model. These results also imply universal generation in the classical sense of Wasserstein GANs arjovsky2017wasserstein.
4.1 Universal Approximation of arbitrary Lipschitz functions between arbitrary dimensions
At first glance, the homeomorphisms and dimensional constraints $d=D$, constraint defining (1), may seem to suggest that our universal flow-based generative models are overly restrictive and thus cannot approximate continuous functions between Euclidean spaces, locally on compact subsets as standard multilayer perceptrons do; cf. hornik1989multilayer; cybenko1989approximation; funahashi1989approximate. However, this is not the case, and in fact the full-power of incremental generative models are not needed if one is prepared to sacrifice injectivity. The idea is similar to Klee’s trick, cf. klee1955some, where we encode transport the graph of an arbitrary continuous function into a homeomorphism. Rather, we can do better, by embedding a function into a single very simple time-$1$ flow up to conjuration be simple linear maps; our simple approach is summarized in Figure 4.
Fix $d,D\in\mathbb{N}_{+}$ and $L\geq 0$. Then, for every $L$-Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ induces a $\bar{L}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\max\{1,L\}$-Lipschitz vector field $V_{f}:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ defined for each $(x,y)\in\mathbb{R}^{d+D}$ by
$V_{f}(x,y)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\big(0,f(x)\big).$ (12)
The solution $\Phi:[0,\infty)\times\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ to the autonomous ODE induced by $V_{f}$ defined by
$\frac{dx_{t}^{(x,y)}}{dt}=0,\qquad\frac{dy_{t}^{(x,y)}}{dt}=f(x_{t}^{(x,y)}),\mbox{ where }(x_{0},y_{0})=(x,y)$
is easily explicitly solved:
$\frac{dx_{t}^{(x,y)}}{dt}=0\Rightarrow x_{t}^{(x,y)}=x\Rightarrow y_{t}^{(x,y)}=y_{0}+\int_{0}^{t}f(x)dt\Rightarrow y_{t}^{(x,y)}=y+t\,f(x)$
and yields the $\bar{L}$-Lipschitz map
$\Phi(t,(x,y))=\big(x,y+t\,f(x)\big).$ (13)
Therefore, the time-$1$ flow of $V_{f}$, induced by $f$, is $\operatorname{Flow}{}(V_{f})(x,y)=(x,y+f(x))$. Restricting this flow to the “lifted hypercube” $0,1]^{d\uparrow}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{(x,y)\in\mathbb{R}^{d+D}:\,x\in[0,1]^{d},\mbox{ and }y=0\}$, which is a compact subset of $0,1]^{d+D}$, we find that for any $(x,0)\in[0,1]^{d\uparrow}$ we have $\operatorname{Flow}{}(V_{f})(x,0)=(x,f(x)).$ Now, let $\pi_{D}^{d+D}:\mathbb{R}^{d+D}\to\mathbb{R}^{D}$ denote the canonical (linear) projection sending any $(x,y)\in\mathbb{R}^{d+D}$ to $y\in\mathbb{R}^{D}$ and let $\iota_{d}^{d+D}:\mathbb{R}^{d}\to\mathbb{R}^{d+D}$ denotes the (linear) embedding sending any $x\in\mathbb{R}^{d}$ to $(x,0)\in\mathbb{R}^{d+D}$; in particular, note that $\iota_{d}^{d+D}([0,1]^{d})=[0,1]^{d\uparrow}$. Putting it all together, we have that: for all $x\in[0,1]^{d}$
$\pi_{D}^{d+D}\circ\operatorname{Flow}{}(V_{f})\circ\iota_{d}^{d+D}(x)=f(x).$ (14)
Applying Theorem 5 to $f$, we conclude that by post- and pre-composing our incremental flow-based models, we can uniformly approximate any continuous function on the cube $0,1]^{d}$.
(a) Lift of: $f(x)=\sin(x)$ to $V(x,y)=(0,\sin(x))$.
(b) Lift of: $f(x)=e^{\tfrac{|x|}{2\pi}}$ to $V(x,y)=(0,e^{\tfrac{|x|}{2\pi}})$.
Figure 4: Any continuous function $f:\mathbb{R}^{d}\to\mathbb{R}$ can be realized as a time-$1$ flow $\Phi(1,(x,y))$ for the $(d+1)$-dimensional vector field $V(x,y)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(0,f(x))$ (illustrated by the pink vector fields), which acts trivially in its first “dummy” coordinates and acts as the target function $f(x)$ in the $(d+1)^{\mathrm{st}}$ coordinate. By mapping any given input $x\in\mathbb{R}^{d}$ to the initial condition $(x,0)\in\mathbb{R}^{d+1}$ of the flow $\Phi$ “lifting” $f$, we then simply flow linearly relative to the $(d+1)^{\mathrm{st}}$ (here $y$) axis until it arrives at $(x,f(x))$ at time 11, at which point the final value can be linearly projected-off and the value $f(x)$ is recovered. In this way, every real-valued continuous (resp. Lipschitz, resp. smooth) function can be realized as a time-$1$ flow of the same regularity in a space of only one more dimension.
However, we note that the model in (14) does not scale well when $D$ is large as one would obtain an approximation rate of $\mathcal{O}(n^{d+D})$ which is significantly slower than the “unconstrained” optimal rate achievable by ReLU MLPs; cf. yarotsky2018optimal. Instead, a rate nearly equal to the “unconstrained” ReLU MLP rate is possible if we alternatively approximate each component of the target function $f$ independently using this strategy, and then concatenate the resulting lifted flow-based approximations. Denoting $f=(f_{1},\dots,f_{D})$, we replace (14) with the model
$\bigoplus_{i=1}^{D}\,\Big(\pi_{1}^{d+1}\circ\operatorname{Flow}{}(V_{f_{i}})\circ\iota_{d}^{d+1}(x)\Big)=\big(f_{1}(x),\dots,f_{D}(x)\big)=f(x).$ (15)
The advantage of the representation in (15), with its greater width, over the more “naive narrow” version in (14), is that each $\operatorname{Flow}{}(V_{f_{i}})$ performs its approximation in only one extra dimension beyond the physical dimension $d$. This design minimizes the approximation-theoretic difficulties that typically arise from high dimensionality. Importantly, doing so achieves the minimax optimal approximation rates (cf. (SHEN2022101, Theorem 2.4) and yarotsky2017error) for the lifted, and thus higher-dimensional space $0,1]^{d+1}$ and nearly achieves the optimal rate on the original low-dimensional domain $0,1]^{d}$, up to an extra factor of $1$; which is possibly inevitable due to the extra invertibility structure of the incremental models studied herein.
Corollary 12 (Approximation of Arbitrary Lipschitz Functions by Linear Lifting)
Let $d,D\in\mathbb{N}_{+}$, $L\geq 0$, and $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ be L-Lipschitz. For every $n\in\mathbb{N}_{+}$ and each $i\in[D]$ there exists ReLU Neural ODEs $\Psi_{i}:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ such that the Latent Neural ODE $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigoplus_{i=1}^{D}(\pi_{1}^{d+1}\circ\Psi_{i}\circ\iota_{d}^{d+1})$ satisfies
$\big\|f(x)-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\in\mathcal{O}\big(\tfrac{1}{n}\big).$
Moreover, $\Psi$ is Lipschitz and the ReLU MLPs parameterizing the vector fields defining each $\Psi_{i}$, have width $\mathcal{O}(dn^{d+1})$ and depth $\mathcal{O}(\log_{2}(d))$ and $\mathcal{O}(dn^{d+1})$ non-zero parameters for each $i\in[D]$.
Thus, Corollary 12 provides a quantitative and, more topologically more explicit, version of the very recent qualitative result of de2025approximation as a direct consequence of our main theorem.
Remark 13 (Optimality of the $\mathcal{O}(n^{d+1})$ rate)
The question of the optimality of the rate $\mathcal{O}(n^{d+1})$ is not known, as there are no available approximation theoretic lower bounds, nor tools for establishing lower bounds, in our non-vector space setting it is currently unknown. However, is likely that a rate of $\mathcal{O}(n^{d})$ cannot be achieved while requiring that the core of the model is the conjugation of a homeomorphism by linear maps, due to the need to lift in order to approximate general continuous functions.
Corollary 14 (Universal Approximation)
Let $d,D\in\mathbb{N}_{+}$. The set $\mathcal{F}\subseteq C([0,1]^{d},\mathbb{R}^{D})$ of all maps of the form (15) where $f:\mathbb{R}^{d}\to\mathbb{R}^{D}$ is a ReLU MLP of arbitrary depth and width is dense in $C([0,1]^{d},\mathbb{R}^{D})$.
Discussion: Only one flow is enough when lifting
If $f$ were smooth, then the topological embedding $x\mapsto(x,f(x))$ defines a differentiable $d$-cell, in the sense of (palais1960extending, page 274). By (palais1960extending, Theorem C), this embedding extends to a compactly supported, orientation-preserving homeomorphism $\Psi:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$, which can then be approximated using our main positive result (Theorem 5). Projecting away the second coordinate, as in (15), yields an approximation of $f$. Moreover, the smoothness assumption is not restrictive: Compactly supported homeomorphisms are identity out of their support, this means that there exists a $r$ such that they are all identity on the boundary of the disc $D^{r}$. By Alexander’s trick, this means they are all isotopic to identity. As a compactly supported homeomorphism of $\mathbb{R}^{d}$ for $d\geq 5$ can be approximated uniformly by compactly supported diffeomorphisms if and only if it is isotopic to a diffeomorphism by  MULLER2014315 this shows in our case working on diffeomorphisms in high dimension is not restrictive. A similar approach was recently explored in pmlr-v162-puthawala22a.
We highlight that this line of reasoning misses the central insight of the flow-based construction in (15). Namely, it does not specify how many flows are required to represent $\Psi$. In contrast, our construction shows that only one flow suffices if we allow for linear lifting/projecting, thereby breaking the homeomorphism structure in (1).
We examine the implications of our results to classical generative modelling, in the sense of Wasserstein GANs arjovsky2017wasserstein; korotin2021wasserstein, and highlight the differences and similarities.
4.2 Universal Approximation of measures satisfying Caffarelli Conditions
We first recall that for any hyperparameter $1\leq p<\infty$, we denote by $\mathcal{P}_{p}([0,1]^{d})$ the subset of probabilities that finitely integrate $x\mapsto\|x\|^{p}$. We equip $\mathcal{P}_{p}([0,1]^{d})$ with the Wasserstein $p$-distance $\mathcal{W}_{p}$, that is, for $\mu,\nu\in\mathcal{P}_{p}([0,1]^{d})$, the metric defined by
$\mathcal{W}_{p}(\mu,\nu)^{p}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\inf_{\pi\in\operatorname{Cpl}(\mu,\nu)}\mathbb{E}_{(X,Y)\sim\pi}\big[\|X-Y\|^{p}\big],$
where $\operatorname{Cpl}(\mu,\nu)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{\pi\in\mathcal{P}([0,1]^{d}\times[0,1]^{d})\colon\pi\text{ has first marginal }\mu,\text{ second marginal }\nu\}$. Given any $n,m\in\mathbb{N}_{+}$, any Borel map $g:[0,1]^{n}\to[0,1]^{m}$, and any probability measure $\nu\in\mathcal{P}([0,1]^{n})$ we recall that the pushforward measure is $g_{\sharp}\nu\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\nu(g^{-1}[\cdot])$ belongs to $\mathcal{P}([0,1]^{m})$. If $g$ is Lipschitz and $\nu$ belongs to the Wasserstein space $\mathcal{P}_{p}([0,1]^{n})$ then so does $g_{\sharp}\nu\in\mathcal{P}_{p}([0,1]^{m})$. We require the following standard regularity condition.
Assumption 4.1 (Caffarelli Conditions; cf. caffarelli1996boundary)
Let $\mu$ and $\nu$ be measures which are supported on $(0,1)^{d}$ and absolutely continuous with respect to the uniform measure $U$ on $(0,1)^{d}$; whose Radon-Nikodym densities $\tfrac{d\mu}{dU}$ and $\tfrac{d\nu}{dU}$ are $\alpha$-Hölder for some $\alpha\in(0,1)$, and are bounded (above and below on $(0,1)^{d}$).
Armed with these definitions we are ready to show that incremental flow-based generation, augmented by lifts, are universal generative models in the more general but less structured context where GANs are typically studied; lu2020universal with related guarantees in biau2020some.
Corollary 15
Let $2<d\in\mathbb{N}_{+}$ and $\mu,\nu$ be probability measures on $0,1]^{d}$ satisfying Assumption 4.1. For every $\varepsilon,\delta>0$, there exists a large enough $N\in\mathbb{N}_{+}$ and a Neural ODE $\Psi:\mathbb{R}^{d+D}\to\mathbb{R}^{d+D}$ such that for all i.i.d. random variables $Z_{1}\sim\dots\sim Z_{N}$, with law $\nu$ (defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$)
$\mathbb{P}\Big(\mathcal{W}_{1}(\mu,\Psi_{\sharp}\nu^{(N)})\leq L^{\Psi}\tfrac{\sqrt{d}C_{[0,1]^{d}}}{\sqrt[d]{N}}+\delta+\varepsilon\Big)\geq 1-2\,e^{\tfrac{-2N\delta^{2}}{d(L^{\Psi})^{2}}}$
where $\nu^{(N)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tfrac{1}{N}\,\sum_{n=1}^{N}\,\delta_{Z_{n}}$ is the empirical distribution induced from the random sample and $C_{[0,1]^{d}}$ is a constant.
5 Conclusion
In this article, we have shown that incremental generation is both necessary and sufficient for universal flow-based approximation, while non-incremental flows form a non-universal subset. By coupling dynamical, algebraic, and approximation-theoretic tools, we established quantitative rates and demonstrated that lifted incremental flows yield structured universality for both functions and probability measures.
Acknowledgments and Disclosure of Funding
The authors would like to thank Paulo Varandas for his encouraging suggestions and for verifying our arguments in proving the meagreness of the set of flows and to Giulia Livieris for her incredibly helpful feedback during the final stages of the manuscript. The authors would also like to thank Andrew Warren for the helpful references.
A. Kratsios and H. Rouhvarzi would like to acknowledge financial support from an NSERC Discovery Grant No. RGPIN-2023-04482 and No. DGECR-2023-00230. They also acknowledge that resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute [^5] https://vectorinstitute.ai/partnerships/current-partners/.
Appendix A Proof of the Negative Result
Proposition 16 (Existence of a recurrent point)
For every compact set $M$ and any $f\in\operatorname{\mathcal{H}}_{d}(M)$ there exists a recurrent point in $M$.
Proof : Take $f\in\operatorname{\mathcal{H}}_{d}(M)$. Since $M$ is compact the set
$\mathcal{A}=\{S\subseteq M:S\neq\emptyset,S\text{ closed},f(S)=S\}$
is non-empty (e.g. $M\in\mathcal{A}$). Order $\mathcal{A}$ by inclusion. Any totally ordered chain $\{S_{i}\}$ in $\mathcal{A}$ has
$\bigcap_{i}S_{i}\;\neq\;\varnothing$
by compactness, and $\bigcap_{i}S_{i}$ is closed and $f$-invariant. By Zorn’s Lemma there is a minimal element $S_{\min}\in\mathcal{A}$, i.e. a nonempty closed invariant set containing no proper nonempty closed invariant subset. Take any $x\in S_{\min}$. If $x$ was not recurrent then there would exist a neighborhood $U\ni x$ and an integer $N$ such that
$f^{n}(x)\notin U\quad\forall\,n\geq N.$
Set
$S^{\prime}\;=\;\overline{\{\,f^{n}(x):n\geq N\}}.$
Then $S^{\prime}$ is nonempty, closed, and $f$-invariant, but $x\notin S^{\prime}$, so $S^{\prime}\subsetneq S_{\min}$, contradicting minimality. Thus $x$ must be recurrent. Since $x$ was arbitrary in $S_{\min}$, every point of $S_{\min}$ is recurrent.
In order to prove Lemma 19 We restate two lemmas for the convenience of the reader. For the proof refer to Lemma 9 and 13 of Nit_1975__FilDecExp :
Lemma 17
Given $\varepsilon>0$ and a flow $\phi$. Suppose $\gamma$ is a $C^{1}$ curve in $M$ (an embedded closed interval or circle) such that at each point $x$ in the image of $\gamma$ one of the following conditions holds:
  1. (i)
    $\|\dot{\phi}(x)\|<\varepsilon/2$, or
  2. (ii)
    $x\notin\mathcal{Z}(\phi)$, and $\gamma$ has inclination $\sigma<\varepsilon/\|\dot{\phi}\|$ at $x$.
Then, given any neighbourhood $U$ of the image of $\gamma$, there exists a flow $\psi$ on $M$ satisfying:
  1. (a)
    $\dot{\psi}=\dot{\phi}$ off $U$,
  2. (b)
    $\|\dot{\psi}-\dot{\phi}\|<\varepsilon$ on $M$,
  3. (c)
    $\gamma$ is (a segment of an) integral curve of $\psi$.
Lemma 18
Let $M$ be a manifold of dimension $\geq 2$ with distance $d$ coming from a Riemannian metric. Suppose a finite collection $\{(p_{i},q_{i})\in M\times M:\,i=1,\ldots,k\}$ of pairs of points of $M$ is specified, together with a small positive constant $\delta>0$ such that:
  1. (a)
    For each $i$, $d(p_{i},q_{i})<\delta$.
  2. (b)
    If $i\neq j$, then $p_{i}\neq p_{j}$ and $q_{i}\neq q_{j}$.
Then there exists $f\in\mathrm{Diff}(M)$ such that
  1. (i)
    $d\bigl(f(x),x\bigr)<2\pi\delta$ for every $x\in M$.
  2. (ii)
    $f(p_{i})=q_{i}$ for $i=1,\ldots,k$.
Lemma 19 ($C^{0}$-closing lemma for compactly supported functions)
Let $f\colon M\to M$ be a compactly supported homeomorphism on the compact set $K\subset M$, where $M$ is a $\mathcal{C}^{\infty}$ manifold of dimension $\geq 2$ with distance $d$ coming from a Riemannian metric, and let $x_{0}\in M$ be a non-wandering point. Then for every $\varepsilon>0$ and compact set $P$ s.t. $K\subset\text{int}(P)$ ($K$ is compactly embedded in $P$) there exists a compactly supported homeomorphism $g\colon M\to M$ on $P$ such that $x_{0}$ is a periodic point of $g$ and
$d_{1}(f,g)<\varepsilon$
also, if $M$ is orientable and $f$ is orientation-preserving, then $g$ is also orientation-preserving.
Proof : The Lemma is trivial when $x_{0}$ is a fixed point or a periodic point. Consider not.
If $M$ is compact and $K=M$ by (AnnaA_1996__C0closing, Theorem 4) the result follows. Suppose not.
As $f$ is compactly supported, $\mathcal{O}_{f}(x_{0})$ is a subset of $K$. Note $f$ is uniformly continuous on $P$ and Lemma 17 and Lemma 18 are true on any Riemannian manifold like $M$ so one can apply the same proof of (AnnaA_1996__C0closing, Theorem 4) which results the desirable function $g=f\circ h$ which $h$ is isotopic to the identity (considering the flow time as the isotopy parameter) which means $h$ is orientation-preserving. As a result, if $f$ is orientation-preserving then so is $g$.
Also, Lemma 17 property (a) makes the perturbed function in Lemma 18 to have support on a finite number of neighborhoods of the curves joining $p_{i}$ to $q_{i}$. As $h$ is the result of Lemma 18 take $\eta$ and those neighbourhoods small enough such that these neighbourhoods stay inside $P$. This makes the function $h$ to be compactly supported on $P$. As a result, $g$ must be compactly supported on $P$.
We are now in a position to prove our main result in topological dynamics which will imply our main negative result (namely Theorem 2).
Proof of Theorem 3:
[…]
$f_{1}(z)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\mathrm{exp}_{x}\circ[\rho\cdot R+(1-\rho)\cdot F]\circ\mathrm{exp}_{x}^{-1}(z)$
[…]
As $D\text{exp}_{x}(0)=Id$ there exists $L>1$ tending to 1 as $r\rightarrow 0$ s.t. :
$d(f_{1}(z),f(z))\leq L\|(F-R)\circ\mathrm{exp}_{x}^{-1}(z)\|\leq L\|F-R\|r\leq L\|F-R\|\varepsilon$ (16)
[…]
For every $C^{0}$-Perturbation $g$ of $f_{2}$ in $\operatorname{\mathcal{H}}_{d}(P)$ one has: $0\neq\text{deg}(f_{2}^{k}(x)-x,0)=\text{deg}(g^{k}(x)-x,0)$. There exists a point $z$ s.t. $g^{k}(z)-z=0$ so $g^{k}$ has a fixed point. Thus, $g$ has a periodic point of period $k>1$ that is not contained in a continuum subset of the set of $k$-periodic points. It means $f_{2}$ has an open neighbourhood. This proves the openness.
Proof of Theorem 2: Since every neural ODE is, by definition, a flowable homeomorphisms supported on any such $\delta,1-\delta]^{d}$ then the set of neural ODEs we mentioned is itself meagre in $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Proof of Theorem 1: Suppose by contradiction that for every $\varphi\in\operatorname{\mathcal{H}}_{d}([0,1]^{d})$ there exists $\varphi_{\sigma}\in\mathrm{NODE}_{\sigma}([0,1]^{d})$ such that $\|\varphi_{\sigma}-\varphi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\frac{\epsilon}{2}$. Then there exists $\delta>0$ and a bump function $\psi$ which is one on $\delta,1-\delta]^{d}$ and zero out of $\frac{\delta}{2},1-\frac{\delta}{2}]^{d}$ and $\|\varphi_{\sigma}\circ\psi-\varphi_{\sigma}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\frac{\epsilon}{2}$; then, $\|\varphi_{\sigma}\circ\psi-\varphi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon$. As $\epsilon$ is arbitrary and $\varphi_{\sigma}\circ\psi\in\mathrm{NODE}_{\sigma}((0,1)^{d})$ this contradicts with Theorem 2.
Appendix B Proof of the Positive Result
First we handle ”Quantitative Formulations” and so we can easily prove the theorem in ”Qualitative Universal Approximation Guarantee” section.
B.1 Quantitative Formulations
B.1.1 General Case
First we prove Proposition 7 to show it is reasonable to think of incremental flow-based generation.
Proof of Proposition 7: Consider vector fields $\tilde{V}(x,y)=(-\pi(y-\frac{1}{2}),\pi(x-\frac{1}{2}))$ and $\tilde{W}(x,y)=(-x+\frac{1}{2},0)$ which are $\pi$ rotation around the point $p=(\frac{1}{2},\frac{1}{2})$ and squeeze to the line $x=\frac{1}{2}$ (call this line $l$). Consider the bump function $\rho$ which is $1$ on the circle with radius $r=\frac{1}{8}$ around $p$ and zero out of the circle with radius $R=\frac{1}{4}$. $V=\tilde{V}\circ\rho,W=\tilde{W}\circ\rho$ are compactly supported on $0,1]^{d}$. Denote time one map of $V,W$ as $\varphi_{V},\varphi_{W}$. Then $\varphi_{V}\circ\varphi_{W}\notin\mathrm{Flow}$ . suppose not, i.e. there exists a vector field $F$ which its time one flow $\varphi_{F}$ is equal to $\varphi_{V}\circ\varphi_{W}$.
Note that for the composition $\varphi_{V}\circ\varphi_{W}$ any point $a\in B_{r}(p)\setminus l$ converges to zero on iteration but any $b\in l$ has periodic point two. Consider the trajectory $\gamma$ of the flow of $\varphi_{V}\circ\varphi_{W}$. For any point $c=r_{c}e^{i\theta}$ (considering polar coordinate) on a neighbourhood $S$ of the point $\varphi_{V}\circ\varphi_{W}(b)$ on $\gamma$, radius of $\varphi_{V}\circ\varphi_{W}(c)$ will be smaller than $r_{c}$ as a result radius of any point other than $b$ on $(\varphi_{V}\circ\varphi_{W})^{2}(S)$ will be smaller but $(\varphi_{V}\circ\varphi_{W})^{2}(b)=b$ so there will be a discontinuity at the point $b$ but non-fixed periodic point of time one flow has to lie on an invariant embedded circle but the only invariant set is the line $l$, which is a contradiction.
Now, we first show how one can approximate a flow, then we consider the composition of flows and prove Theorem 8.
Lemma 20 (Transfer: Universal Approximation of Vector Field to Flow)
Let $n\in\mathbb{N}$ and $V\in\mathcal{X}^{0}$, then for every $\varphi=\operatorname{Flow}{\left({V}\right)}$ there exists a ReLU MLP $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ compactly supported on $(0,1)^{d}$ with width $8d(n+1)^{d}$, depth $\lceil\text{log}_{2}d\rceil+6$, and at-most $16d(n+1)^{d}+9$ non-zero parameters such that the flow (Neural ODE) given for each $x\in\mathbb{R}^{d}$ by
$\Psi(x)$ $=z_{1}^{x}$ (17)
$z_{t}^{x}$ $=x+\int_{0}^{t}\,\Phi(z_{s}^{x})\,ds$
for $0\leq t\leq 1$; satisfies the uniform estimates
$\|\Psi(x)-\varphi(x)\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(\frac{d}{2n})\|_{l^{\infty}}e^{L^{V}}$ (18)
Where $\omega$ is the modulus of regularity of $\varphi$ which is equal to the modulus of regularity of $\Phi$. Note that $e^{L^{V}}$ does not depend on $n$. One gets convergence as $n\rightarrow\infty$
Proof : If we define $(V)_{j}:=V_{j}$ then as $V_{j}$ is Lipschitz and nonzero on a compact subset from (Hong_2024__ReLUMLPs) there exists a ReLU MLP $(\phi(z_{s}^{x}))_{j}:=\phi_{j}(z_{s}^{x})$ and modulus of regularity $(\omega)_{j}:=\omega_{j}$ such that:
$\|V_{j}(z_{s}^{x})-\phi_{j}(z_{s}^{x})\|_{L^{\infty}([0,1]^{d})}\leq\omega_{j}(\frac{d}{2n})$
To make it compactly supported on $\frac{\delta}{4},1-\frac{\delta}{4}]^{d}$ for a $\delta\in(0,2)$ we compose it componentwise with ReLU bump function below:
$\begin{split}b(x)&=\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{\delta}{2})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{\delta}{2}))\bigr)\\ &=\begin{cases}0,&\text{if }x<\frac{\delta}{4}\\ 2x-\frac{\delta}{2}&\text{if }x\in[\frac{\delta}{4},\frac{\delta}{2}]\\ x,&\text{if }x\in[\frac{\delta}{2},1-\frac{\delta}{2}]\\ (1-\frac{1}{\delta})x+(\frac{2-\delta}{2\delta}),&\text{if }x\in[1-\frac{\delta}{2},1-\frac{\delta}{4}]\\ 0,&\text{if }x>1-\frac{\delta}{4}\end{cases}\end{split}$
implemented by the following network:
$\mathbf{x}$ $\Rightarrow\begin{bmatrix}\sigma_{ReLU}(x-\frac{\delta}{4})\\ \sigma_{ReLU}(x-\frac{2\delta}{4})\\ \sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\end{bmatrix}$
$\Rightarrow\bigl[\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{2\delta}{4})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\bigr)]$
$\Rightarrow\bigl[\phi_{j}\bigl(\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{2\delta}{4})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\bigr)\bigr)]=\phi_{j}\!\bigl(b(x)\bigr)$
$\Rightarrow\Phi_{j}(x)$
which is $2$-Lipschitz and supported on $[0,1]$. The result $\Phi_{j}=\phi_{j}\circ b$ has width:
$\mathrm{width}(\Phi_{j})$ $=\mathrm{max}\{3d,d,\mathrm{width}(\phi_{j})\}$
$=\mathrm{max}\{3d,d,8d(n+1)^{d}\}$
$=8d(n+1)^{d}$
$=\mathrm{width}(\phi_{j})$
And depth:
$\mathrm{depth}(\phi_{j}\circ b)=\mathrm{depth}(\phi_{j})+2=\lceil log_{2}d\rceil+6$
then note that $\Phi_{j}=\phi_{j}$ on $\frac{\delta}{2},1-\frac{\delta}{2}]^{d}$ so:
$\|\Phi_{j}-V_{j}\|_{L^{\infty}([\frac{\delta}{2},1-\frac{\delta}{2}]^{d})}\leq\|\phi_{j}-V_{j}\|_{L^{\infty}([0,1]^{d})}\leq\omega_{j}(\frac{d}{2n})$
And:
$\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,\frac{\delta}{2}]^{d})}\leq\frac{\delta}{2}(\|V_{j}\|_{L^{\infty}([0,1]^{d})}+\|\phi_{j}\|_{L^{\infty}([0,1]^{d})})$
Then:
$\begin{split}\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,1]^{d})}&\leq\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,\frac{\delta}{2}]^{d})}+\|\Phi_{j}-V_{j}\|_{L^{\infty}([\frac{\delta}{2},1-\frac{\delta}{2}]^{d})}+\|\Phi_{j}-V_{j}\|_{L^{\infty}([1-\frac{\delta}{2},1]^{d})}\\ &\leq\omega_{j}(\frac{d}{2n})+\delta(\|V_{j}\|_{L^{\infty}([0,1]^{d})}+\|\phi_{j}\|_{L^{\infty}([0,1]^{d})})\end{split}$
Note that $\mathrm{supp}(\Phi)=\mathrm{supp}(V)=[0,1]^{d}$ so we can consider the the inequalities above on $\mathbb{R}^{d}$ then we’ll have:
$\begin{split}\|V(x_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&\leq\|V(x_{s}^{x})-V(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|V(z_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq L^{V}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\omega(\frac{d}{2n})\|_{l^{\infty}}+\delta(\|V\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\phi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}})\end{split}$
Note that as $\phi$ and $V$ are continuous on a compact set, they attain their supremum. This means there exists a constant $C$ such that:
$\|V\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\phi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq C$
Observe that, $C\geq 0$ not only depends on $V$ but also on $n$ and on $d$, as $\phi$ depends only on $V$ and $n$ and on $d$; hence by Jensen’s inequality:
$\begin{split}\|z_{t}^{x}-x_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\|\int_{0}^{t}V(x_{s}^{x})-\Phi(z_{s}^{x})ds\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\int_{0}^{t}\|V(x_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds\\ &\leq L^{V}\int_{0}^{t}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds+\left(\|\omega(\frac{d}{2n})\|_{l^{\infty}}+\delta C\right)t\end{split}$
On the other hand, observe that $\delta$ does not depend on any of $V,n,d$ and it was arbitrary; thus, we may take $\delta C\leq\|\omega(\frac{d}{2n})\|_{l^{\infty}}$ thus:
$\begin{split}\|z_{t}^{x}-x_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq L^{V}\int_{0}^{t}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds+2\|\omega(\frac{d}{2n})\|_{l^{\infty}}t\end{split}$
By applying Grönwall’s inequality (Pachpatte1998 Theorem 1.3.1)
$\|z_{t}^{x}-x_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(\frac{d}{2n})\|_{l^{\infty}}te^{L^{V}t}$
let $t=1$ :
$\|\Psi(x)-\varphi(x)\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}=\|z_{1}^{x}-x_{1}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(\frac{d}{2n})\|_{l^{\infty}}e^{L^{V}}$ (19)
The right hand side goes to zero as $n\rightarrow\infty$. Finally, observe that $e^{L^{V}}$ depends only on $V$ and not on the approximation parameter $n\in\mathbb{N}_{+}$.
Now we are ready to prove the Theorem 8.
Proof of  Theorem 8: As $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ in (7) define $\varphi_{t}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{t})$ for every $t\in[T]$. So one can write:
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)$ (20)
We use Lemma 20 for each $\varphi^{(t)}$ to find Neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}$. One can write $\Psi^{(t)}(x)=z_{1}^{x}$ as in (17), then as $\Phi^{(t)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(t)}(z_{s}^{x})-\Phi^{(t)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi^{(t)}}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
By applying Grönwall’s inequality (Pachpatte1998 Theorem 1.2.2)
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{(t)}}t}$
Let $t=1$ :
$\|\Psi^{(t)}(x)-\Psi^{(t)}(y)\|_{l^{\infty}}=\|z_{1}^{x}-z_{1}^{y}\|_{l^{\infty}}\leq\|x-y\|_{l^{\infty}}e^{L^{\Phi^{(t)}}}$ (21)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\Psi^{(t)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ using induction. The base case is already proved in inequality (21); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|\leq\small\prod_{t=1}^{T-1}e^{L^{{\Phi^{{(t)}}}}}\big\|x-y\big\|$
Using (21) we can write:
$\big\|\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(y)\big\|\leq e^{L^{\Phi^{(T)}}}\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|\leq\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|$ (22)
So one can deduce the Lipschitz constant is at-most $\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$.
We again use induction to show inequality (9). The base case is already proved in Lemma 20. For the step of the induction, consider the induction assumption bellow:
$\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T-1}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T-1}e^{L^{V^{(j)}}}\right)$
Using (29) we can write:
$\begin{split}\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{t=1}^{T}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)-\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{(T)}}}+2\|\omega^{(T)}(\frac{d}{2n})\|_{l^{\infty}}e^{L^{V^{(T)}}}\end{split}$
Where for the last inequality we used (18) and (21).
Then by the induction assumption:
$\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(\frac{d}{2n})\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$
By Lemma 20 the right hand side goes to zero as $n\rightarrow\infty$. We have thus concluded the proof.
B.1.2 Differentiable Case
We do the same steps for the differentiable case but with another ReLU MLP which gives faster approximation rates. Also we first prove a proposition which Proposition 9 is a special case of it.
Lemma 21 (Transfer: Universal Approximation of Vector Field to Flow (Differentiable Case))
Let $s\in\mathbb{N}_{+}$, and $V=(V_{1},\dots,V_{d})\in\mathcal{X}^{s}$, and $\omega_{j}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}{85(s+1)^{d}8^{s}\|V_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}}$ for $1\leq j\leq d$ and $\omega\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\omega_{1},\dots,\omega_{d})$ then for every $\varphi=\operatorname{Flow}{\left({V}\right)}$ and for any $N_{j},L\in\mathbb{N}^{+}$ , $1\leq j\leq d$, There exists a ReLU MLP
$\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
compactly supported in $(0,1)^{d}$ with width less than or equal $\sum_{j=1}^{d}17sd^{\,d+13}d(N_{j}+2)\log_{2}(8N_{j})$ and depth $18s^{2}(L+2)\log_{2}(4L)+2(d+1)$ such that the flow (Neural ODE) given for each $x\in\mathbb{R}^{d}$ by
$\Psi(x)$ $=z_{1}^{x}$ (23)
$z_{t}^{x}$ $=x+\int_{0}^{t}\,\Phi(z_{s}^{x})\,ds$
for $0\leq t\leq 1$; satisfies the uniform estimates
$\|\Psi(x)-\varphi(x)\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}e^{L^{V}}$ (24)
where $\Psi$ is also compactly supported in $(0,1)^{d}$.
Proof : If we define $(V)_{j}:=V_{j}$ for $1\leq j\leq d$ then by  (lu2021deep, Theorem 1.1) for any $N_{j},L\in\mathbb{N}^{+}$, there exists a function $\phi_{j}$ implemented by a ReLU FNN with width $C_{1}(N_{j}+2)\log_{2}(8N_{j})$ and depth $C_{2}(L+2)\log_{2}(4L)+2d$ such that
$\|\phi_{j}-V_{j}\|_{L^{\infty}([0,1]^{d})}\leq C_{3}\|V_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d},$ (25)
where $C_{1}=17s^{d+1}3^{d}d$, $C_{2}=18s^{2}$, and $C_{3}=85(s+1)^{d}8^{s}$.
Let $\delta\in(0,2)$, one can make $\phi$ compactly supported on $\frac{\delta}{4},1-\frac{\delta}{4}]^{d}$ by composing it componentwise with ReLU bump function below:
$\begin{split}b(x)&=\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{\delta}{2})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{\delta}{2}))\bigr)\\ &=\begin{cases}0,&\text{if }x<\frac{\delta}{4}\\ 2x-\frac{\delta}{2}&\text{if }x\in[\frac{\delta}{4},\frac{\delta}{2}]\\ x,&\text{if }x\in[\frac{\delta}{2},1-\frac{\delta}{2}]\\ (1-\frac{1}{\delta})x+(\frac{2-\delta}{2\delta}),&\text{if }x\in[1-\frac{\delta}{2},1-\frac{\delta}{4}]\\ 0,&\text{if }x>1-\frac{\delta}{4}\end{cases}\end{split}$
implemented by the following network:
$\mathbf{x}$ $\Rightarrow\begin{bmatrix}\sigma_{ReLU}(x-\frac{\delta}{4})\\ \sigma_{ReLU}(x-\frac{2\delta}{4})\\ \sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\end{bmatrix}$
$\Rightarrow\bigl[\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{2\delta}{4})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\bigr)]$
$\Rightarrow\bigl[\phi_{j}\bigl(\sigma_{ReLU}\bigl(2\sigma_{ReLU}(x-\frac{\delta}{4})-\sigma_{ReLU}\!(x-\frac{2\delta}{4})-\frac{1}{\delta}\sigma_{ReLU}\!(x-(1-\frac{2\delta}{4}))\bigr)\bigr)]=\phi_{j}\!\bigl(b(x)\bigr)$
$\Rightarrow\Phi_{j}(x)$
which is $2$-Lipschitz and supported on $[0,1]$. The result $\Phi_{j}=\phi_{j}\circ b$ has width:
$\mathrm{width}(\Phi_{j})$ $=\mathrm{max}\{3d,d,\mathrm{width}(\phi_{j})\}$
$=\mathrm{max}\{3d,d,17sd^{\,d+13}d(N_{j}+2)\log_{2}(8N_{j})\}$
$=17sd^{\,d+13}d(N_{j}+2)\log_{2}(8N_{j})$
$=\mathrm{width}(\phi_{j})$
And depth:
$\mathrm{depth}(\phi_{j}\circ b)=\mathrm{depth}(\phi_{j})+2=18s^{2}(L+2)\log_{2}(4L)+2(d+1)$
and as $\mathrm{depth}(\phi_{1})=\mathrm{depth}(\phi_{2})=\dots=\mathrm{depth}(\phi_{d})$ then $\mathrm{depth}(\Phi_{1})=\mathrm{depth}(\Phi_{2})=\dots=\mathrm{depth}(\Phi_{d})$, by parallelization  (petersen2024mathematical, Proposition 2.3.) the neural network
$\Phi(x)=(\Phi_{1}(x),\ldots,\Phi_{d}(x)):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
has the same depth. The width of it is at least :
$\mathrm{width}(\Phi)\leq\sum_{j=1}^{d}\mathrm{width}(\Phi_{j})$
Set $\omega_{j}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}{C_{3}\|V_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}}$ , $\phi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\phi_{1},\dots,\phi_{d})$ and $\omega\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}(\omega_{1},\dots,\omega_{d})$ then note that $\Phi_{j}=\phi_{j}$ on $\frac{\delta}{2},1-\frac{\delta}{2}]^{d}$ so:
$\|\Phi_{j}-V_{j}\|_{L^{\infty}([\frac{\delta}{2},1-\frac{\delta}{2}]^{d})}\leq\|\phi_{j}-V_{j}\|_{L^{\infty}([0,1]^{d})}\leq C_{3}\|V_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d},$
Also on the interval $0,\frac{\delta}{2}]^{d}$ one can write:
$\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,\frac{\delta}{2}]^{d})}\leq\frac{\delta}{2}(\|V_{j}\|_{L^{\infty}([0,1]^{d})}+\|\phi_{j}\|_{L^{\infty}([0,1]^{d})})$
Then:
$\begin{split}\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,1]^{d})}&\leq\|\Phi_{j}-V_{j}\|_{L^{\infty}([0,\frac{\delta}{2}]^{d})}+\|\Phi_{j}-V_{j}\|_{L^{\infty}([\frac{\delta}{2},1-\frac{\delta}{2}]^{d})}+\|\Phi_{j}-V_{j}\|_{L^{\infty}([1-\frac{\delta}{2},1]^{d})}\\ &\leq C_{3}\|V_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}+\delta(\|V_{j}\|_{L^{\infty}([0,1]^{d})}+\|\phi_{j}\|_{L^{\infty}([0,1]^{d})})\end{split}$
Note that $\mathrm{supp}(\Phi)=\mathrm{supp}(V)=[0,1]^{d}$ so we can consider the the inequalities above on $\mathbb{R}^{d}$ then we’ll have:
$\begin{split}\|V(x_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&\leq\|V(x_{s}^{x})-V(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|V(z_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq L^{V}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\\ &\quad\quad+\delta(\|V\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\phi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}})\end{split}$
Note that as $\phi$ and $V$ are continuous on a compact set, they attain their supremum. This means there exists a constant $C$ such that:
$\|V\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}+\|\phi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq C.$
Observe that, $C\geq 0$ not only depends on $V$ but also on $N_{1},\dots,N_{d}$ and on $L$, as $\phi$ depends only on $V$ and $N_{1},\dots,N_{d}$ and on $L$; hence by Jensen’s inequality:
$\begin{split}\|x_{t}^{x}-z_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\|\int_{0}^{t}V(x_{s}^{x})-\Phi(z_{s}^{x})ds\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\int_{0}^{t}\|V(x_{s}^{x})-\Phi(z_{s}^{x})\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds\\ &\leq L^{V}\int_{0}^{t}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds+\left(\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}+\delta C\right)t\end{split}$ (26)
On the other hand, observe that $\delta$ does not depend on any of $V,N_{1},\dots,N_{d},L$ and it was arbitrary; thus, we may take $\delta C\leq\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}$ thus:
$\|x_{t}^{x}-z_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq L^{V}\int_{0}^{t}\|x_{s}^{x}-z_{s}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}ds+2\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}t$ (27)
By applying Grönwall’s inequality (Pachpatte1998 Theorem 1.3.1)
$\|z_{t}^{x}-x_{t}^{x}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}te^{L^{V}t}$
Let $t=1$ :
$\|\Psi(x)-\varphi(y)\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}=\|z_{1}^{x}-x_{1}^{y}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq 2\|\omega(N_{1},\dots,N_{d},L)\|_{l^{\infty}}e^{L^{V}}.$
Proposition 22 (Universal Approximation by Deep Neural ODEs (Differentiable Case))
Let $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ and $1\leq s$; then, there exists ReLU neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ satisfies the approximation guarantee:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\end{split}$ (28)
Where $(\omega^{(t)})_{j}=85(s+1)^{d}8^{s}\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}({N_{j}}{L})^{-2s/d}$ each $\omega^{(t)}$. The right hand side goes to zero as $L\rightarrow\infty$ and $N_{j}\rightarrow\infty$ for each $j\in[T]$.
Moreover, $\Psi$ is a compactly-supported homeomorphism t on $\mathbb{R}^{d}$ supported on $(0,1)^{d}$, with Lipschitz constant at-most $\prod_{t=1}^{T}e^{L^{\Psi^{(t)}}}$ and $\Psi^{(t)}$ is a ReLU MLP of width less than or equal to $\sum_{j=1}^{d}17s^{d+1}3^{d}d(N^{t}_{j}+2)\log_{2}(8N^{t}_{j})$ and depth $18s^{2}(L^{t}+2)\log_{2}(4L^{t})+2(d+1)$ where $N^{t}_{j},L^{t}\in\mathbb{N}^{+}$ for each $1\leq j\leq d,$ and $1\leq t\leq T$.
In particular, $L^{\Psi}$ depends on the parameters $N_{1},\dots,N_{d}$.
Proof : As $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$ in (7) define $\varphi_{t}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\operatorname{Flow}{}(V_{t})$ for every $t\in[T]$. So one can write:
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)$ (29)
We use Lemma 21 for each $\varphi^{(t)}$ to find Neural ODEs $\{\Psi^{(t)}\}_{t=1}^{T}$. One can write $\Psi^{(t)}(x)=z_{1}^{x}$ as in (23), then as $\Phi^{(t)}$ is Lipschitz we can write:
$\big\|z_{t}^{x}-z_{t}^{y}\big\|_{l^{\infty}}\leq\big\|x-y\big\|_{l^{\infty}}+\int_{0}^{t}\big\|\Phi^{(t)}(z_{s}^{x})-\Phi^{(t)}(z_{s}^{y})\big\|_{l^{\infty}}ds\leq\big\|x-y\big\|_{l^{\infty}}+L^{\Phi^{(t)}}\int_{0}^{t}\big\|z_{s}^{x}-z_{s}^{y}\big\|_{l^{\infty}}ds$
By applying Grönwall’s inequality (Pachpatte1998 Theorem 1.2.2)
$\|z_{t}^{x}-z_{t}^{y}\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(t)}}}t}$
set $t=1$:
$\big\|\Psi^{(t)}(x)-\Psi^{(t)}(y)\big\|_{l^{\infty}}=\big\|z_{1}^{x}-z_{1}^{y}\big\|_{l^{\infty}}\leq\left(\|x-y\|_{l^{\infty}}\right)e^{L^{\Phi^{{(t)}}}}$ (30)
Define $\Psi\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\Psi^{(t)}$. We try to show that the Lipschitz constant $L^{\Psi}$ is at-most $\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$ using induction. The base case is already proved in inequality (30); for the step of the induction, consider the induction assumption bellow :
$\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq\small\prod_{t=1}^{T-1}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|_{l^{\infty}}$
Using (30) we can write:
$\big\|\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq e^{L^{\Phi^{(T)}}}\big\|\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(y)\big\|_{l^{\infty}}\leq\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}\big\|x-y\big\|_{l^{\infty}}$ (31)
So one can deduce the Lipschitz constant is at-most $\small\prod_{t=1}^{T}e^{L^{\Phi^{(t)}}}$.
Note that one can apply the same procedure of (30) for $\varphi^{(t)}$ and conclude
$L^{\varphi^{(t)}}=e^{L^{V^{(t)}}}$ (32)
We again use induction to show inequality (28). The base case is already proved in Lemma 21. For the step of the induction consider the induction assumption bellow:
$\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T-1}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T-1}e^{L^{V^{(j)}}}\right)$
Using (29) we can write:
$\begin{split}\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&=\big\|\bigcirc_{t=1}^{T}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T}\,\varphi^{(t)}(x)-\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\quad\quad+\big\|\varphi^{(T)}\circ\bigcirc_{t=1}^{T-1}\Psi^{(t)}(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\\ &\leq\big\|\bigcirc_{t=1}^{T-1}\varphi^{(t)}(x)-\bigcirc_{t=1}^{T-1}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}e^{L^{V^{{(T)}}}}\\ &\quad\quad+2\|\omega^{(T)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}e^{L^{V^{{(T)}}}}\end{split}$
Where for the last inequality we used (24) and (32).
Then by the induction assumption:
$\big\|\varphi(x)-\bigcirc_{t=1}^{T}\,\Psi^{(t)}(x)\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N_{1},\dots,N_{d},L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)$
By Lemma 21 the right hand side goes to zero as $L\rightarrow\infty$ and $N_{j}\rightarrow\infty$ for each $j\in[d]$. We have thus concluded the proof.
Now we are in a position to prove Proposition 9.
Proof of Proposition 9: Take free parameters $N=N_{1}=N_{2}=\dots=N_{d}$ then the result follows from Proposition 22
B.1.3 Smooth Case
First we prove Theorem 10 and then Proposition 11.
Lemma 23 (Normality)
The set $\operatorname{Flow}{}$ generates a normal subgroup (denoted as) $\langle\operatorname{Flow}{}\rangle$ of $\operatorname{Diff}_{0}([0,1]^{d})$.
Proof : If $\varphi=\operatorname{Flow}{}(V)$ then using Corollary 9.14 (Lee_2013__SmoothManifolds) in our case it shows that for every $F\in\operatorname{Diff}_{0}([0,1]^{d})$ the Flow of the (pushforward) vector field $F_{*}V$ is $\eta=F\circ\varphi\circ F^{-1}$, in other words $\eta=\operatorname{Flow}{}(F_{*}D)\Rightarrow\eta\in\operatorname{Flow}{}$ so for $\varphi$ in the generated subgroup $\langle\operatorname{Flow}{}\rangle$ and the diffeomorphism $F\in\operatorname{Diff}_{0}([0,1]^{d})$ we can write :
$\varphi\in\langle\operatorname{Flow}{}\rangle\Rightarrow\exists n\in\mathbb{N}_{+}\,\text{ s.t. }\forall i\in\{1,\dots,n\},\;\exists\varphi^{(i)}\in\operatorname{Flow}{}(V_{i})\text{ s.t. }\varphi=\varphi^{(n)}\circ\dots\circ\varphi^{(1)}$
then one can write:
$F\circ\varphi\circ F^{-1}=(F\circ\varphi^{(n)}\circ F^{-1})\circ(F\circ\varphi^{(n-1)}\circ F^{-1})\circ\dots\circ(F\circ\varphi^{(1)}\circ F^{-1})=\eta^{(n)}\circ\dots\circ\eta^{(1)}\in\langle\operatorname{Flow}{}\rangle$
Lemma 24 (Large incrementality Triviality for Smooth Regularity)
For $\delta>0$, there exists $T_{d}\in\mathbb{N}_{+}$ s.t. for any integer $T>T_{d}$, $\operatorname{\mathcal{H}}_{d}^{T,\infty}((-\delta,1+\delta)^{d})$ is empty.
Proof : Define $S:=\{\,\operatorname{Flow}{}(V)\mid V\text{ smooth, compactly supported in }(0,1)^{d}\,\}$ and the norm
$\|f\|_{\mathrm{frag}}=\begin{cases}0,&f=\mathrm{id},\\[6.0pt] \displaystyle\min\{\,m\in\mathbb{N}\mid f=h_{1}\cdots h_{m},\;h_{i}\in S\,\},&f\neq\mathrm{id}.\end{cases}$
[…]
Lemma 25 (Decomposition)
For every diffeomorphism $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$, there exist some $T\in\mathbb{N}$ and flows $\varphi^{(1)},\dots,\varphi^{(T)}\in\operatorname{Flow}{}$ such that
$\varphi=\bigcirc_{t=1}^{T}\,\varphi^{(i)}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\varphi^{(T)}\circ\dots\circ\varphi^{(1)}.$ (33)
Proof : By Thurston’s Theorem, see e.g. (Banyaga_1997__DiffeoBook, Theorem 2.1.1), the group $\operatorname{Diff}_{0}([0,1]^{d})$ is simple; i.e. it has no proper normal subgroups besides the trivial group consisting only of the identity (diffeomorphism) on $\mathbb{R}^{d}$. Since $\operatorname{Flow}{}$ does not contain only the identity diffeomorphism on $\mathbb{R}^{d}$ and, by Lemma (23), since $\operatorname{Flow}{}$ generates a normal subgroup of $\operatorname{Diff}_{0}([0,1]^{d})$ then $\operatorname{Flow}{}$ generates the entire group $\operatorname{Diff}_{0}([0,1]^{d})$. Consequentially, each $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ admits a representation of the form (33).
Remark 26
above theorems can be applied on any $a,b]^{d},\ (a,b)^{d}$ for $a<b\in\mathbb{R}$.
Now we are in a position to prove Theorem 10.
Proof of Theorem 10: By Lemma 25 there exists a $T$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})$; as $\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})$ is non-empty and for $0<\delta$ we have $\operatorname{\mathcal{H}}_{d}^{T,\infty}([0,1]^{d})\subset\operatorname{\mathcal{H}}_{d}^{T,\infty}((-\delta,1+\delta)^{d})$, by applying Lemma 24 we conclude $T\leq T_{d}$. This means one can write any diffeomorphism as a composition of at most $T_{d}$ flows.
Using Theorem 10 we can prove Proposition 11.
Proof of Proposition 11: As $\varphi\in\operatorname{Diff}_{0}([0,1]^{d})$ it is also $s$ differentiable so by Theorem 10 there exists $T\leq T_{d}$ such that $\varphi\in\operatorname{\mathcal{H}}_{d}^{T,s}([0,1]^{d})$. By Proposition 9 there exists $\Psi$ such that:
$\begin{split}\big\|\varphi-\Psi\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}&\leq\sum_{t=1}^{T}\left(2\|\omega^{(t)}(N,L)\|_{l^{\infty}}\prod_{j=t}^{T}e^{L^{V^{(j)}}}\right)\\ &\leq 170(s+1)^{d}8^{s}\max_{t,j}\{\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}\}(NL)^{-2s/d}e^{\max_{t}\{L^{V^{(t)}}\}}T_{d}\left(\prod_{j=2}^{T}L^{V^{(j)}}\right)\\ &\in O\big(N^{-2s/d}\big)\end{split}$ (34)
as $(\omega^{(t)})_{j}=85(s+1)^{d}8^{s}\|V^{(t)}_{j}\|_{C^{s}([0,1]^{d})}(NL)^{-2s/d}$
B.2 Qualitative Universal Approximation Guarantee
We are now able to deduce our main qualitative universal approximation guarantee within the class $\operatorname{\mathcal{H}}_{d}([0,1]^{d})$.
Proof of Theorem 5: Now, since $\varphi$ is supported on $B(0,R)$ for large enough $R$ then there exists some $M>0$ such that $\varphi$ is supported on $M,M]^{d}\supset B(0,R)$. Recalling by a contraction/expansion diffeomorphism $\varphi_{M}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ of the form $\varphi_{M}(x)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}Ax$ for some $d\times d$ scalar matrix $A=kI_{d}$, for some $k>0$, satisfying $\varphi_{M}([-M,M])\subseteq B(0,1/8)$. Moreover $\varphi_{M}$ can be written as the time $1$ flow whose integral curve with initial condition $x\in\mathbb{R}^{d}$ is
$x_{t}^{x}=k^{t}x\mbox{ for all }t\geq 0$
We may without loss of generality consider $\mathrm{supp}^{\dagger}(\varphi\circ\varphi_{M}^{-1})=B(0,\frac{1}{8})$ thus $\varphi\circ\varphi_{M}^{-1}$ fixes all points on a neighbourhood of the boundary of $B(0,1/4)$. Consequently, we may apply the Munkres-Connel-Bing Theorem, as formulated in (muller2014uniform, Lemma 2) to deduce that: for every $\varepsilon>0$ there exists a diffeomorphism $\varphi_{\varepsilon}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ supported on $B(0,\frac{1}{4})$ satisfying the uniform approximation guarantees
$\|\varphi_{\varepsilon}(x)-\varphi\circ\varphi_{M}^{-1}(x)\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\tfrac{\varepsilon}{3}.$ (35)
Which by substituting $x=\varphi_{M}(x)$ we’ll still have:
$\|\varphi_{\varepsilon}\circ\varphi_{M}-\varphi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\tfrac{\varepsilon}{3}.$ (36)
Now, since $\varphi_{\varepsilon}$ is supported on $B(0,\frac{1}{4})\subset[0,1]^{d}$ and since $\varphi_{\varepsilon}$ is a diffeomorphism then Theorem 10 implies that there exists some $\tilde{T}_{d}$ (not depending on $\varphi_{\varepsilon}$ nor on $\varepsilon$) and some $1\leq T\leq\tilde{T}_{d}$ such that $\varphi_{\varepsilon}\in\operatorname{\mathcal{H}}_{d}^{T,1}([0,1]^{d})$. Applying Theorem 8 implies that there exists ReLU neural ODEs $\{\Psi^{(t)}=\mathrm{Flow}(\Phi^{(t)})\}_{t=1}^{T}\subset\mathcal{X}^{0}$ such that the homeomorphism $\Psi_{\varepsilon}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\bigcirc_{t=1}^{T}\,\Psi^{(t)}$ is Lipschitz and compactly supported and satisfies the approximation guarantee:
$\big\|\varphi_{\epsilon}-\Psi_{\varepsilon}\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\tfrac{\varepsilon}{3}.$ (37)
As $\varphi_{M}\in\operatorname{\mathcal{H}}_{d}([-M,M]^{d})$ by the same theorem:
$\big\|\varphi_{M}-\Psi_{M}\big\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}\leq\tfrac{\varepsilon}{3L^{\Psi_{\varepsilon}}}.$ (38)
Combining (36), (37), (38):
$\begin{split}\big\|\Psi_{\varepsilon}\circ\Psi_{M}-\varphi\big\|&\leq\big\|\Psi_{\varepsilon}\circ\Psi_{M}-\varphi_{\epsilon}\circ\varphi_{M}\big\|+\big\|\varphi_{\epsilon}\circ\varphi_{M}-\varphi\big\|\\ &\leq\big\|\Psi_{\varepsilon}\circ\Psi_{M}-\Psi_{\varepsilon}\circ\varphi_{M}\big\|+\big\|\Psi_{\varepsilon}\circ\varphi_{M}-\varphi_{\epsilon}\circ\varphi_{M}\big\|+\big\|\varphi_{\epsilon}\circ\varphi_{M}-\varphi\big\|\\ &\leq L^{\Psi_{\varepsilon}}\big\|\Psi_{M}-\varphi_{M}\big\|+\big\|\Psi_{\varepsilon}\circ\varphi_{M}-\varphi_{\epsilon}\circ\varphi_{M}\big\|+\big\|\varphi_{\epsilon}\circ\varphi_{M}-\varphi\big\|\\ &\leq L^{\Psi_{\varepsilon}}\tfrac{\varepsilon}{3L^{\Psi_{\varepsilon}}}+\tfrac{\varepsilon}{3}+\tfrac{\varepsilon}{3}=\varepsilon\end{split}$ (39)
and setting $\Psi=\Psi_{\varepsilon}\circ\Psi_{M}$ and $T_{d}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{T}_{d}+1$ yields the conclusion. [^9] Note that because $\varphi_{M}\in\operatorname{\mathcal{H}}_{d}([-M,M]^{d})$ we have $\mathrm{supp}^{\dagger}(\Psi)=[-M,M]^{d}$
Appendix C Proofs of Corollaries
Proof of Corollary 12: By Theorem 8 and using (15), there exists Lipschtiz ReLU neural ODEs $\{\Psi_{i}=\operatorname{Flow}{}(\Phi_{i})\}_{i=1}^{D}$ such that for each $i\in[D]$ we have:
$\|\operatorname{Flow}{}(V_{f_{i}})-\Psi_{i}\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon$
Which by Theorem 8, $\varepsilon\in O\big(\tfrac{1}{n}\big)$ so restricting the domain and only considering the last coordinate of the output we’ll have:
$\|\pi_{1}^{d+1}\circ\operatorname{Flow}{}(V_{f_{i}})\circ\iota_{d}^{d+1}-\pi_{1}^{d+1}\circ\Psi_{i}\circ\iota_{d}^{d+1}\|_{L^{\infty}(\mathbb{R}^{d})}<\epsilon$
but by (14) we know $f_{i}=\pi_{1}^{d+1}\circ\operatorname{Flow}{}(V_{f_{i}})\circ\iota_{d}^{d+1}$ so define $\Psi=\bigoplus_{i=1}^{D}(\pi_{1}^{d+1}\circ\Psi_{i}\circ\iota_{d}^{d+1})$ then:
$\|f-\Psi\|_{L^{\infty}(\mathbb{R}^{d})l^{\infty}}<\epsilon.$
also width and depth are given in table 1. Also note that by lifting, projecting and concatenation the function remains Lipschitz. This completes our proof.
Proof of Corollary 14: The Benyamini-Lindenstrauss theorem; see e.g. (BenyaminiLindenstrauss_2000_NonlinearFunctionalAnalysis, Theorem 1.12), implies that for every non-empty subset $A\subseteq[0,1]^{d}$, each $L\geq 0$, and every $L$-Lipschitz function $g:A\to\mathbb{R}^{D}$, there exists a $L$-Lipschitz extension $G:[0,1]^{d}\to\mathbb{R}^{D}$; i.e. $g|_{A}=G$. Therefore, (Miculescu_LipDense, Theorem 1) implies that the set of Lipschitz functions is dense in $C([0,1]^{d},\mathbb{R}^{D})$ for the uniform topology. The result now follows from Corollary 12.
Proof of Corollary 15:
Applying the Benamou-Brenier Theorem (VillaniOTTopics2003, Theorem 2.12 (ii) and (iii)), we deduce that there exists a convex function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ inducing the unique optimal transport map between $\mu$ and $\nu$, namely $\mu=\nabla\varphi_{\sharp}\nu$ i.e. $\varphi$ is a Kantorovich potential. note that, at this state $\nabla\varphi$ may only be defined $\mu$-a.s. therefore we verify its regularity before being able to continue further with any uniform approximation thereof. Under Assumption 4.1 the Caffarelli’s regularity theorem, as formulated (VillaniOT2009, Theorem 12.50 (ii)) applies and we deduce that $\varphi\in C^{2,\alpha}((0,1)^{d})$; in particular, $\nabla\varphi$ is defined on all of $(0,1)^{d}$ and it’s Lipschitz thereon by the mean-valued theorem.
Since $\nabla\varphi:(0,1)^{d}\to(0,1)^{d}$ is $L^{\prime}$-Lipschitz, for some $L^{\prime}\geq 0$, then it admits an $L^{\prime}$-Lipschitz extension $\tilde{\varphi}:[0,1]^{d}\to[0,1]^{d}$, again by the Benyamini-Lindenstrauss Theorem, see e.g. (BenyaminiLindenstrauss_2000_NonlinearFunctionalAnalysis, Theorem 1.12) (it is easy to see that it must map the close cube to itself, since $\nabla\varphi$ maps $(0,1)^{d}$ to itself and the latter is continuous).
Fix $\varepsilon>0$, applying Corollary 12 we find that there exists a Lipschitz Latent Neural ODE $\Psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ induced by ReLU MLP of depth $\mathcal{O}(\log_{2}(d))$, width $\mathcal{O}(dn^{d+1})$, with $\mathcal{O}(dn^{d+1})$ non-zero parameters, such that
$\Big(\max_{x\in[0,1]^{d}}\,\big\|\tilde{\varphi}(x)-\Psi(x)\big\|^{2}\Big)^{1/2}<\varepsilon$ (40)
Since $\mu$ and $\nu$ are both supported on $0,1]^{d}$ then they belong to $\mathcal{P}_{p}([0,1]^{d})$ for every $1\leq p<\infty$; in particular, $\mu$ and $\Psi_{\sharp}\nu$ belong to $\mathcal{P}_{2}(\mathbb{R}^{d})$. Therefore, $\mathcal{W}_{2}(\mu,\Psi_{\sharp}\nu)$ is finite and we have
$\mathcal{W}_{1}\big(\mu,\Psi_{\sharp}\nu^{(N)}\big)\leq\mathcal{W}_{1}\big(\Psi_{\sharp}\nu,\Psi_{\sharp}\nu^{(N)}\big)+\mathcal{W}_{1}\big(\mu,\Psi_{\sharp}\nu\big)\leq\underbrace{\mathcal{W}_{1}\big(\Psi_{\sharp}\nu,\Psi_{\sharp}\nu^{(N)}\big)}_{\text{(I)}}+\underbrace{\mathcal{W}_{2}\big(\mu,\Psi_{\sharp}\nu\big)}_{\text{(II)}}.$ (41)
We begin by controlling term (I) . Indeed, since $\Psi$ is $L^{\Psi}$-Lipschitz, then we have
$\text{(I)}=\mathcal{W}_{1}\big(\Psi_{\sharp}\nu,\Psi_{\sharp}\nu^{(N)}\big)\leq\operatorname{Lip}\big(\Psi\big)\,\mathcal{W}_{1}\big(\nu,\nu^{(N)}\big).$ (42)
Now, applying the concentration of measure result in  (hou2023instance, Lemma 18) we have that: for every $\frac{\delta}{L^{\Psi}}>0$
$\mathbb{P}\biggl(\Big|\mathcal{W}_{1}\big(\nu,\nu^{(N)}\big)-\mathbb{E}\big[\mathcal{W}_{1}\big(\nu,\nu^{(N)}\big)\big]\Big|\geq\frac{\delta}{L^{\Psi}}\biggr)\leq 2\,e^{\tfrac{-2N\delta^{2}}{d(L^{\Psi})^{2}}}\,\,\mbox{ and }\,\,\mathbb{E}\big[\mathcal{W}_{1}\big(\nu,\nu^{(N)}\big)\big]\leq\tfrac{\sqrt{d}C_{[0,1]^{d}}}{\sqrt[d]{N}}$
for some constant $C_{[0,1]^{d}}>0$. Consequently, we may control (I) probabilistically: for every $\delta>0$
$\mathbb{P}\Big(L^{\Psi}\mathcal{W}_{1}\big(\nu,\nu^{(N)}\big)\leq L^{\Psi}\tfrac{\sqrt{d}C_{[0,1]^{d}}}{\sqrt[d]{N}}+\delta\Big)\geq 1-2\,e^{\tfrac{-2N\delta^{2}}{d(L^{\Psi})^{2}}}$ (43)
So
$\mathbb{P}\Big(\text{(I)}\leq L^{\Psi}\tfrac{\sqrt{d}C_{[0,1]^{d}}}{\sqrt[d]{N}}+\delta\Big)\geq 1-2\,e^{\tfrac{-2N\delta^{2}}{d(L^{\Psi})^{2}}}$ (44)
It remains to control term (II). In particular,
$\text{(II)}=\mathcal{W}_{2}\big(\mu,\Psi_{\sharp}\nu\big)=\mathcal{W}_{2}\big(\nabla\varphi_{\sharp}\nu,\Psi_{\sharp}\nu\big),$ (45)
Since $0,1]^{d}\setminus(0,1)^{d}$ is of $\nu$-measure zero, we may again rephrase (45) as
$\text{(II)}=\mathcal{W}_{2}\big(\nabla\varphi_{\sharp}\nu,\Psi_{\sharp}\nu\big)\leq\mathbb{E}_{X\sim\nu}\big[\|(\nabla\varphi)_{\sharp}(X)-(\Psi)_{\sharp}(X)\|^{2}\big]^{1/2}=\mathbb{E}_{X\sim\nu}\big[\|\tilde{\varphi}(X)-(\Psi)_{\sharp}(X)\|^{2}\big]^{1/2}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\text{(III)}.$ (46)
But
$\text{(II)}\leq\text{(III)}\leq\Big(\max_{x\in[0,1]^{d}}\,\big\|\tilde{\varphi}(x)-\Psi(x)\big\|^{2}\Big)^{1/2}$ (47)
Consequently, combining (40), (44) ,and (47) yields the conclusion.
Appendix D Additional Geometric and Topological Background
  1. 1.
    continuum: a compact and connected metric space that contains at least two points.
  2. 2.
    Tangent vector, tangent bundle and vector field: Let $M$ be a smooth $n$-manifold.
    1. (a)
      A tangent vector at a point $p\in M$ can be defined as the velocity $\gamma^{\prime}(0)$ of a smooth curve $\gamma:(-\varepsilon,\varepsilon)\to M$ with $\gamma(0)=p$, or equivalently as a derivation at $p$, i.e. a linear map $v:C^{\infty}(M)\to\mathbb{R}$ satisfying the Leibniz rule
      $v(fg)=v(f)g(p)+f(p)v(g),\qquad f,g\in C^{\infty}(M).$
      The set of all tangent vectors at $p$ forms an $n$-dimensional vector space called the tangent space at $p$, denoted $T_{p}M$.
    2. (b)
      The tangent bundle of $M$ is the disjoint union
      $TM=\bigsqcup_{p\in M}T_{p}M,$
      together with the natural projection $\pi:TM\to M$ given by $\pi(v)=p$ for $v\in T_{p}M$. It is itself a smooth $2n$-dimensional manifold.
    3. (c)
      A vector field on $M$ is a smooth map
      $X:M\to TM$
      such that $X(p)\in T_{p}M$ for every $p\in M$. Equivalently, a vector field is a smooth section of the tangent bundle $\pi:TM\to M$. In local coordinates $(x^{1},\dots,x^{n})$, any vector field has the form
      $X=\sum_{i=1}^{n}f_{i}(x)\,\frac{\partial}{\partial x^{i}},$
      where $f_{i}\in C^{\infty}(M)$.
  3. 3.
    Homotopic and $C^{k}$-diffeotopy: A homotopy between two continuous functions $f$ and $g$ from a topological space $X$ to a topological space $Y$ is defined to be a continuous function
    $H:X\times[0,1]\to Y$
    such that $H(x,0)=f(x)$ and $H(x,1)=g(x)$ for all $x\in X$.
    Let $M$ and $N$ be smooth manifolds, and let $f_{0},f_{1}:M\to N$ be $C^{k}$ diffeomorphisms ($1\leq k\leq\infty$).
    We say that $f_{0}$ and $f_{1}$ are $C^{k}$-diffeotopic if there exists a map
    $F:M\times[0,1]\to N$
    such that:
    1. (a)
      For each $t\in[0,1]$, the map
      $f_{t}(x):=F(x,t):M\to N$
      is a $C^{k}$ diffeomorphism.
    2. (b)
      $f_{0}=F(\cdot,0)$ and $f_{1}=F(\cdot,1)$.
    3. (c)
      The map $F$ is $C^{k}$ in $x$ and continuous (sometimes $C^{k}$) in $(x,t)$.
  4. 4.
    Conjugation invariant and autonomous norm: conjugation-invariant norm $\nu:G\to[0;+\infty)$ on a group $G$ is a function which satisfies the following axioms:
    • (i)
      $\nu(1)=0;$
    • (ii)
      $\nu(f)=\nu(f^{-1})\quad\forall f\in G;$
    • (iii)
      $\nu(fg)\leq\nu(f)+\nu(g)\quad\forall f,g\in G;$
    • (iv)
      $\nu(f)=\nu(gfg^{-1})\quad\forall f,g\in G;$
    • (v)
      $\nu(f)>0$ for all $f\neq 1$
    for a smooth function $f$ define $\|f\|_{\mathrm{frag}}:=\min\left\{m\in\mathbb{N}\mid f=h_{1}\cdots h_{m}\text{ where }h_{i}=\mathrm{Flow}(V_{i})\text{ for some }V_{i}\in\mathcal{X}\right\}$ where $\text{Flow}(V_{i})$ is the time one solution of Cauchy Problem for smooth vector field $V_{i}$
  5. 5.
    Nowhere dense and meagreness: Let $T=(S,\tau)$ be a topological space and $\subseteq S. AA$ is nowhere dense in $T$ if and only if
    $\bigl(\overline{A}\bigr)^{\circ}\;=\;\varnothing,$
    where $\overline{A}$ denotes the closure of $A$ and $A^{\circ}$ its interior.
    AA is meagre in $T$ if and only if it can be written as a countable union
    $A=\bigcup_{n=1}^{\infty}A_{n},$
    where each $A_{n}\subseteq S$ is nowhere dense in $T$. Equivalently, complement of $A$ is open and dense.
  6. 6.
    Complete vector field: Let $M$ be a smooth manifold and let $X$ be a smooth vector field. Denote by $\varphi_{t}$ the local flow of $X$, that is, the solution to the ODE
    $\frac{d}{dt}\varphi_{t}(p)=X(\varphi_{t}(p)),\qquad\varphi_{0}(p)=p.$
    We say that $X$ is complete if for every $p\in M$, the integral curve $\gamma_{p}(t):=\varphi_{t}(p)$ is defined for all $t\in\mathbb{R}$. Equivalently, $X$ is complete if its local flow extends to a global flow
    $\varphi:\mathbb{R}\times M\to M.$
  7. 7.
    Portable manifold: We say that a smooth connected open manifold $M$ is portable if it admits a complete vector field $X$ and a compact subset $M_{0}$ with the following properties:
    • $M_{0}$ is an attractor of the flow $X^{t}$ generated by $X$: for every compact subset $K\subset M$ there exists $\tau>0$ so that $X^{\tau}(K)\subset M_{0}$.
    • There exists a diffeomorphism $\theta\in\mathrm{Diff}_{0}(M)$ so that $\theta(M_{0})\cap M_{0}=\emptyset$.
  8. 8.
    Recurrent point: Let $X$ be a topological space and let $f:X\to X$ be a continuous map. A point $x\in X$ is called a recurrent point of $f$ if there exists a sequence of integers $(n_{k})$ with $n_{k}\to\infty$ such that
    $f^{n_{k}}(x)\longrightarrow x\quad\text{as }k\to\infty.$
    Equivalently, $x$ is recurrent if it belongs to its own $\omega$-limit set,
    $x\in\omega(x):=\{y\in X:f^{n_{k}}(x)\to y\ \text{for some sequence }n_{k}\to\infty\}.$
  9. 9.
    Orbit, fixed point and periodic point: Let $X$ be a set and $f:X\to X$ a map. For $x\in X$, the forward orbit of $x$ under $f$ is the set
    $\mathcal{O}^{+}(x):=\{f^{n}(x):n\in\mathbb{N}_{0}\},$
    where $\mathbb{N}_{0}=\{0,1,2,\dots\}$ and $f^{0}=\mathrm{id}_{X}$.
    If $f$ is invertible, the (full) orbit of $x$ is
    $\mathcal{O}(x):=\{f^{n}(x):n\in\mathbb{Z}\}.$
    A point $x\in X$ is called a fixed point of $f$ if
    $f(x)=x.$
    A point $x\in X$ is called a periodic point of period $k\geq 1$ if
    $f^{k}(x)=x,$
    and $k$ is the smallest positive integer with this property.
    The set of all periodic points of period $k$ is denoted
    $\operatorname{Per}_{k}(f):=\{\,x\in X:f^{k}(x)=x\ \text{and}\ f^{j}(x)\neq x\ \text{for all }0<j<k\,\}.$